### v1

# A Fast Algorithms for Fourier extension based on boundary interval data

Abstract
This paper present a new algorithm for the computation of Fourier extension based on boundary data, which can obtain a super-algebraic convergent Fourier approximation for non-periodic functions. The algorithm calculates the extension part through boundary data and connects it with the original function to form a periodic smooth function. By testing the key parameters involved, their impact on the algorithm is clarified and the optimization setting scheme of the parameters is proposed. Compared with FFT, the algorithm only needs to increase the computational complexity by a fixed small amount.
keywords:
Fourier Series, Fourier Extensions, Fourier Continuation
{AMS}
42A10, 65T40, 65T50
1 Introduction
Fourier series approximation for smooth periodic functions is popular because it can offer spectral convergence and be computed numerically via the FFT. However, the situation is changed completely once periodicity is lost. A jump discontinuity is imposed in the approximation at the domain boundaries due to the well known Gibbs phenomenon. Many ways have been developed to ameliorate the ill effects of the Gibbs phenomenon, including the methods based on Gegenbauer polynomials [9, 10, 16], the methods utilizing Padé approximations [7, 8] and the Fourier extension (or continuation) methods [4, 11, 12, 14, 13, 15, 6, 17].
The idea of the existing discrete Fourier extension can be described as follows: Let $f$ be a non-periodic, smooth and sufficiently differentiable function over $[-1,1]$, which is sampled in a predefined set of points $\{t_{\ell}\}\subset[-1,1]$, the discrete Fourier extension $C(f)$ of $f$ is defined by
$C(f)=\sum_{k=-N}^{N}{\bf c}_{k}e^{\text{i}\frac{k\pi}{T}t},\quad T>1,$ (1)
where the coefficients $\{{\bf c}_{k}\}_{k=-N}^{N}$ are chosen to minimize the error
$\min\sum_{\ell}\left|f(t_{\ell})-\sum_{k=-N}^{N}{\bf c}_{k}e^{\text{i}\frac{k\pi}{T}t_{\ell}}\right|^{2}.$ (2)
To obtain a solution to the minimization problem (2), a severely ill-conditioned least-squares system need to be solved. In [4], Boyd has introduced the truncated singular value decomposition (TSVD) to stabilize the calculation process, the accuracy almost up to machine precision. A convergence analysis by Adcock et. al states that for analytic functions the method is at least superalgebraically convergent [2]. Utilizing randomized algorithms, Lyon in [12] provided a fast algorithm for the computation with $T=2$. And in [14], Matthysen et. al presented fast algorithms for arbitrary $T$, exploiting it as a specific variant of bandlimited extrapolation. This idea was further developed in the literature [15] to form a fast algorithm for function approximation on 2D domains.
It should be noted that the goal of the extension process is to effectively connect the data at both ends of the function. We construct the following two functions:
$f_{1}(t)=\left\{\begin{aligned} &-1,&x<-\frac{1}{2}\\ &\tanh\left(\tan\pi t\right),&-\frac{1}{2}\leq t\leq\frac{1}{2}\\ &1,&t>\frac{1}{2}\end{aligned}\right.$ (3)
$f_{2}(t)=\left\{\begin{aligned} &-1,&t<-\frac{1}{2}\\ &\tanh\left(100*\tan\pi t\right),&-\frac{1}{2}\leq t\leq\frac{1}{2}\\ &1,&t>\frac{1}{2}\end{aligned}\right.$
Since the data of the two functions near the endpoints are the same, the same connection data can be used on the interval $[-T,1)$ and $(1,T]$. An illustration is shown in Fig. 1. This raises an issue: does the data inside the interval really help the calculation of the extension? In fact, A very effective method, known as FC-Gram method has been developed earlier, which uses only boundary data (projecting it into a basis of orthogonal polynomials) for extension calculations [5]. Recently, this method has also been extended to the approximation of two-dimensional functions[6]. The numerical implementation of the FC-gram algorithm requires a high-precision environment and it is not superalgebraic convergent.
Figure 1: Periodic extensions of $f_{1}(t)$ and $f_{2}(t)$.
This paper proposes a new fast algorithm for the calculation of Fourier extension, which has the following traits:
  • Only a few discretization points in the boundary interval are used to calculate the extension part, so the method can be computed with FFT speed. Unlike the FC-Gram algorithm, the method in this paper does not need to introduce polynomials.
  • Since the extension part only uses boundary data for calculation, the influence of the singular points inside the function on the approximation is greatly weakened. Therefore, for functions with smooth boundaries and internal oscillations, the proposed algorithm has a smaller re-solution constant than the original full-data algorithm.
All the results in this paper are based on numerical experiments, and we leave the theoretical proof as a topic for future work. The proposed algorithm has good performance and gives clear conclusions on the settings of various parameters. Therefore, we feel that it is of substantial interest for those who use FEs in practice.
1.1 Overview of the paper
In Section 2, we outline the algorithm and explain its structure. The specific calculation method of the extension part based on boundary data is given in Section 3. Section 4 systematically tests the key parameters, points out their influencing factors, and gives the optimization setting scheme. In Section 5, we give the numerical test of the final algorithm and compare it with two other fast algorithms.
2 The outline of the boundary interval algorithm
In this paper, we restrict ourselves to the case where the sampling points are uniform, that is,
$t_{\ell}=\frac{\ell}{M},\quad\ell=-M,\ldots,M.$ (4)
Let the positive integers $m_{\Delta}$ be the number of nodes taken in the boundary interval, and $S_{l},S_{r}$ be the set of corresponding subscripts,
$S_{l}=\{-M,-M+1,\ldots,-M+m_{\Delta}-1\},\quad S_{r}=\{M-m_{\Delta}+1,M-m_{\Delta}+2,\ldots,M\}.$ (5)
Let $T_{\Delta}$ be a positive interger,
$L_{\Delta}=2\times T_{\Delta}\times m_{\Delta},\quad h=\frac{2\pi}{L_{\Delta}},\quad x_{j}=(j-1)h,\quad j=1,2,\ldots,L_{\Delta}.$ (6)
Then let
$J_{\Delta,1}=\{1,2,\ldots,m_{\Delta}\},~{}J_{\Delta,2}=\{\frac{L_{\Delta}}{2}+1,\frac{L_{\Delta}}{2}+2,\ldots,\frac{L_{\Delta}}{2}+m_{\Delta}\},~{}J_{\Delta}=J_{\Delta,1}\cup J_{\Delta,2}.$ (7)
For $j\in J_{\Delta}$, take
$g(x_{j})=\left\{\begin{aligned} &f(t_{M-m_{\Delta}+j}),&j\in J_{\Delta,1},\\ &f(t_{-M+j-\frac{L}{2}-1}),&j\in J_{\Delta,2}.\end{aligned}\right.$ (8)
Then we calculate $g_{c}(x_{j}),j=1,2,\ldots,L_{\Delta}$ according to the method in section 3. Finally, we will obtain a function $f_{c}$ with a period of $2+\lambda$, where
$\lambda=\frac{L_{\Delta}-2m_{\Delta}+2}{2M},$ (9)
and the data of the extension function $f_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\frac{L_{\Delta}}{2}-m_{\Delta}$ in one period $[-1,1+\lambda)$ is
$f_{c}(t_{\ell})=\left\{\begin{aligned} &f(t_{\ell}),&|\ell|\leq M,\\ &g_{c}(x_{m_{\Delta}+\ell-M+1}),&\ell>M.\end{aligned}\right.$ (10)
Fig. 2 shows the basic process of the boundary interval algorithm using the function $f(x)=x^{2}$ as an example. Here, in order to make the graph look clearer, we use a relatively small number of nodes. Figure 1 (a) shows the given data (indicated by the black dotted line) and the boundary data selected for the calculation extension (indicated by the red circle). Figure 1 (b) shows the $g_{c}(x_{j})$ calculated from the boundary data. The data represented by the blue circle is the subsequent extension data required. Figure 1 (c) shows the data of $f_{c}$ in one period formed by combining the original data with the calculated extension data. It can be seen from the calculation process that the calculation of the extension part only depends on the boundary interval data.
(a) Given data and selected boundary data
(b) $g_{c}(x_{j})$ with $T_{\Delta}=4$
(c) $f_{c}(t_{\ell})$
Figure 2: Calculation process of a periodic extension of $f(x)=x^{2}$.
After obtaining the data in one period, the subsequent algorithm can be implemented through FFT. Therefore, whether the algorithm can achieve satisfactory efficiency depends on the computational complexity of the extension part, that is, the size of the parameters $T_{\Delta}$ and $m_{\Delta}$. We will give their selection schemes through testing in Section 4.
3 The calculation of $g_{c}$
Including appropriate normalisation, we let
$\phi_{k}(x)=\frac{1}{\sqrt{L_{\Delta}}}e^{\text{i}kx},$ (11)
Assume $g(x_{j})$ is given by (8), then the Fourier extension $g_{c}(x)$ is defined to be
$g_{c}(x)=\sum_{k=-n_{\Delta}}^{n_{\Delta}}{\bf c}_{k}\phi_{k}(x).$ (12)
where the coefficients ${\bf c}_{k}$ are determined by
$\min_{({\bf c}_{k})}\sum_{j\in J_{\Delta}}\left|\sum_{k=-n_{\Delta}}^{n_{\Delta}}{\bf c}_{k}\phi(x_{j})-g(x_{j})\right|^{2}.$ (13)
If we let
[…]
$l=1,2,\ldots,2m_{\Delta},k=-n_{\Delta},\ldots,n_{\Delta}.$
Then the solution of (13) can be obtained by solving the rectangular system
$A{\bf c}\approx b$ (15)
in a least squares sense. If we let
$N_{\Delta}=\min\{2m_{\Delta},2n_{\Delta}+1\},$ (16)
then the SVD of $A$ can be given as
$A=\sum_{i=1}^{N_{\Delta}}{\bf u}_{i}\sigma_{i}{\bf v}^{T}_{i},$ (17)
where $\sigma_{i}$ are the singular values of $A$ such that
$\sigma_{1}\geq\ldots\geq\sigma_{N_{\Delta}}\geq 0,$ (18)
while ${\bf u}_{i}$ and ${\bf v}_{i}$ are the left and right singular vectors of $A$, respectively. We define
$\gamma=\frac{m_{\Delta}}{n_{\Delta}}.$ (19)
Fig. 3 shows the singular values of $A$ calculated by SVD techniques for various $m_{\Delta}$, $n_{\Delta}$ and $T_{\Delta}$ in MATLAB. It can be seen that the first approximately $\mathcal{O}(n_{\Delta}/T_{\Delta})$ are close to $1$ and then the singular values rapidly decay until the true singular values are masked by machine precision errors.
(a) $n_{\Delta}=100$, $T_{\Delta}=6$
(b) $m_{\Delta}=100$, $T_{\Delta}=20$
(c) $m_{\Delta}=100$, $n_{\Delta}=100$
Figure 3: The singular values of $A$ for various $m_{\Delta}$, $n_{\Delta}$ and $T_{\Delta}$.
The situation of singular values reflects the ill-posedness of (15), regularization needs to be introduced. Referring to the treatment methods in [2], the truncated SVD leas-squares solution to (15) can be given as
${\bf c}_{\tau}=\sum_{\sigma_{i}>\tau}\frac{{\bf u}_{i}^{T}b}{\sigma_{i}}{\bf v}_{i},$ (20)
where the parameter $\tau$ is usually chosen close to the machine precision.
4 Numerical investigation and determination of key parameters
The proposed algorithm involves multiple parameters, and their selection has a substantial impact on the performance of the algorithm. This section will test and analyze them and then provide an optimized scheme.
4.1 Testing the parameter $T_{\Delta}$
Previous research results show that the extension length has a substantial impact on the convergence rate and the resolution power. Generally speaking, a larger extension length can achieve a higher convergence rate but also means a larger resolution constant. For details, refer to [11, 1, 3]. In [1], the authors also point out that the theoretical results may not be witnessed in computations.
For the proposed algorithm, we are more concerned about the influence of $T_{\Delta}$ on the convergence rate, and whether there is any relationship between it and the number of boundary points $m_{\Delta}$, the ratio coefficient $\gamma$, and the frequency of the approximated function. Now we take
$f(t)=\exp(i\pi\omega t)$
to test different parameters. In Fig 4, we give results for the approximation error against $T_{\Delta}$ with different parameters. Here and what follows, the approximation error evaluated as the maximum pointwise error over an equi-sampled grid ten times denser than the one used for construction and the parameter $\tau=1e-14$ (we also tested with $\tau=1e-12,1e-13,1e-15$ and obtained results similar to those shown here.)
(a) $\gamma=1$, $M=500$, $m_{\Delta}=100$.
(b) $\gamma=1$, $M=1000$, $\omega=20$.
(c) $\gamma=2$, $M=500$, $m_{\Delta}=150$.
(d) $\gamma=2$, $M=1000$, $\omega=20$.
(e) $\gamma=1/2$, $M=500$, $m_{\Delta}=50$.
(f) $\gamma=1/2$, $M=1000$, $\omega=20$.
Figure 4: The approximation error against $T_{\Delta}$ for various parameters
Three distinct regions are visible in all figures:
  • A region $I_{1}:=\{T_{\Delta}:1<T_{\Delta}<\hat{T}_{1}\}$, where the error decreases rapidly to near the machine accuracy as $T_{\Delta}$ increases.
  • A region $I_{2}:=\{T_{\Delta}:\hat{T}_{1}<T_{\Delta}<\hat{T}_{2}\}$, where the error is maintained close to the machine accuracy.
  • A region $I_{3}:=\{T_{\Delta}>\hat{T}_{2}\}$, where the error rebounds, and the increase of $T_{\Delta}$ will cause the error to increase instead.
It can be seen that the value of $\hat{T}_{1}$ is influenced by $\gamma$, but not by $\omega$, $m_{\Delta}$, and $M$, while $\hat{T}_{2}$ is influenced by $\gamma$, $M$, and $\omega$. For $\hat{T}_{2}$, we can give an estimate:
$\hat{T}_{2}\sim\frac{M}{\gamma\omega}.$ (21)
For several different values of $\gamma$, we give the approximate values of $\hat{T}_{1}$ (We tested the functions of $\omega$ from $1$ to $50$, and took the mean of the corresponding $T_{\Delta}$ when their error was less than $135e-$ 13 for the first time as the estimate of $\hat{T}_{1}$), which are listed in Table 1.
When the number of nodes $M$ is not large enough, the regions $I_{1}$ and $I_{2}$ are merged into one region. At this time, the error reaches the optimal value near $\hat{T}_{2}$. This phenomenon can be observed from Fig. 5.
Table 1: The approximation value of $\hat{T}_{1}$ for various $\gamma$.
$\gamma$ $8$ $4$ $2$ $1$ $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
$\hat{T}_{1}$ 1.09 1.20 2.3 5.9 12.5 24.6 48.9
(a) $\gamma=1$, $M=100$, $m_{\Delta}=50$.
(b) $\gamma=2$, $M=400$, $m_{\Delta}=100$.
Figure 5: The approximation error against $T_{\Delta}$ for $\hat{T}_{2}<\hat{T}_{1}$.
What will happen if the $T_{\Delta}$ values selected is less than $\hat{T}_{1}$? Figure 6 shows this. We can see that the error will not decrease after reaching a certain threshold, which cannot be improved by increasing the numbers of $M$ and $m_{\Delta}$. It should be noted that this situation also applies to the Fourier extension algorithm using full data in previous literature. In Figure 7, we show the results of the algorithm using full data. It can be seen that when the value of $T$ is smaller than the corresponding $\hat{T}_{1}$, the method cannot achieve machine accuracy. Similarly, this cannot be improved by increasing the number of $M$.
(a) $\omega=20$, $\gamma=1$, $m_{\Delta}=50$.
(b) $\omega=20$, $\gamma=1$, $M=500$.
Figure 6: The approximation error against $T_{\Delta}$ for ${T}_{\Delta}<\hat{T}_{1}$ (If $m_{\Delta}>M$, we take $m_{\Delta}=M$).
(a) $\omega=20$, $\gamma=1$.
(b) $\omega=20$, $\gamma=1/2$.
Figure 7: The approximation error against $M$ for full data FE with $T<\hat{T}_{1}$.
From the above test results, we can see that the selection of $T_{\Delta}$ is crucial to the approximation accuracy. Only when $T_{\Delta}$ is within the interval $[\hat{T}_{1},\hat{T}_{2}]$, the approximation can obtain satisfactory results. The length and position of this interval are affected by the parameter $\gamma$ and the frequency $\omega$ of the approximated function.
4.2 Testing the parameter $m_{\Delta}$
The size of parameter $m_{\Delta}$ determines whether this study is meaningful. This subsection tests the influences of other parameters on it. First, we take $M>{T}_{\Delta}\gamma\omega$ for observation. As can be seen from Fig. 8, as the value of $m_{\Delta}$ increases, the error decreases rapidly and reaches the machine precision after a certain value $\hat{m}_{\Delta}$, while the value of $\hat{m}_{\Delta}$ is independent of $M$ and $\omega$. In Fig. 9, we further show how the error changes with $m_{\Delta}$ for different values of $\gamma$ and $T_{\Delta}$. It can be seen that the value of $\hat{m}_{\Delta}$ decreases with the increase of $T_{\Delta}$. When $\gamma\leq 1$, the value of $\hat{m}_{\Delta}$ at the corresponding $\hat{T}_{1}$ is very close. When $\gamma>1$, the value of $\hat{m}_{\Delta}$ at the corresponding $\hat{T}_{1}$ increases rapidly with the increase of $\gamma$. From the perspective of use, we are more concerned about the $\hat{m}_{\Delta}$ value near $\hat{T}_{1}$. We tested this and the results are listed in Table 2.
Table 2: The approximation values of $\hat{m}_{\Delta}$ for various $\gamma$ and $T_{\Delta}$.
$\gamma$ $8$ $4$ $2$ $1$ $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
$T_{\Delta}$ $1.1$ $1.2$ $2.3$ $6$ $12$ $24$ $50$
$\hat{m}_{\Delta}$ 1820 480 58 17 18 18 19
(a) $M=1000$, $\gamma=2$, $T_{\Delta}=6$.
(b) $\omega=20$, $\gamma=2$, $T_{\Delta}=6$.
Figure 8: The approximation error against $m_{\Delta}$ for various $\omega$ and $M$.
(a) $M=2000$, $\gamma=8$, $\omega=20$.
(b) $M=1000$, $\gamma=4$, $\omega=20$.
(c) $M=500$, $\gamma=2$, $\omega=20$.
(d) $M=500$, $\gamma=1$, $\omega=20$.
(e) $M=500$, $\gamma=1/2$, $\omega=20$.
(f) $M=500$, $\gamma=1/4$, $\omega=20$.
Figure 9: The approximation error against $m_{\Delta}$ for various $\gamma$ and $T_{\Delta}$.
(a) $f(t)=\cos\left(\frac{100}{1+25t^{2}}\right)$
(b) approximation error
Figure 10: The function and its approximation errors against $m_{\Delta}$ for $M=500$, $\gamma=1$, $T_{\Delta}=6$.
It is also worth noting that for functions with internal oscillations and smooth boundaries, an increase in the number of $m_{\Delta}$ will cause the value of $\hat{T}_{2}$ to decrease. At this time, if the value of $M$ is not large enough, the approximation result using a larger $m_{\Delta}$ may be worse. This can be seen from the approximation results for the function $f(t)=\cos\left(\frac{100}{1+25t^{2}}\right)$ in Figure 10.
4.3 Computational complexity and parameter determination
The computational complexity of this algorithm consists of three parts:
  • Calculation of $c_{\tau}$. The computational effort in this step is mainly concentrated on the singular value decomposition of the matrix $A$, which is related to $m_{\Delta}$ and $\gamma$, which can be estimated as:
    • If $\gamma<1$, the computational complexity is
      $\mathcal{O}\left(m_{\Delta}^{2}n_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma}m_{\Delta}^{3}\right)$ (22)
    • If $\gamma\geq 1$, the computational complexity is
      $\mathcal{O}\left(m_{\Delta}n^{2}_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma^{2}}m_{\Delta}^{3}\right)$ (23)
  • Calculation of $g_{c}$. This step can be achieved by IFFT, the computational complexity is
    $O\left(L_{\Delta}\log(L_{\Delta})\right)$ (24)
  • Calculation of Fourier coefficients of $f_{c}$. It can be achieved by FFT, the computational complexity is
    $O\left(M\log(M)\right).$ (25)
Combining the test results in the previous two subsections, we know that if we want to approximate the function with frequency $\omega$ with machine precision, the parameters $m_{\Delta}$ and $T_{\Delta}$ need to satisfy
$m_{\Delta}\geq\hat{m}_{\Delta},\quad T_{\Delta}\geq\hat{T}_{1}.$ (26)
Accordingly, combined with formula (21), the parameters $L_{\Delta}$ and $M$ need to satisfy
$L_{\Delta}\geq\hat{L}_{\Delta}=:2\times\hat{T}_{1}\times\hat{m}_{\Delta},\quad M\geq\hat{M}=:\hat{T}_{1}\gamma\omega.$ (27)
Referring to the results in Table 1 and Table 2, the value of $\hat{M}$ is very close for different $\gamma$ and the value of $\hat{L}_{\Delta}$ is relatively close when $\gamma=1$ and $2$, but the value of $\hat{L}_{\Delta}$ is much larger when $\gamma$ takes other values. Further combined with the value of $\hat{m}_{\Delta}$, it can be seen that the computational complexity of the method is equivalent when $\gamma=1$ and $\gamma=2$, while the computational complexity increases significantly in other cases. Furthermore, note that a larger $m_{\Delta}$ means a greater probability of encountering high frequencies in the selected boundary interval. Therefore, we recommend using the following parameters
$\gamma=1,T_{\Delta}=6,m_{\Delta}=20,$ (28)
for actual calculations, and parameters
$\gamma=2,T_{\Delta}=2.3,m_{\Delta}=60,$ (29)
can be used as an alternative. Under such parameter settings, the size of the matrix $A$ in equation (15) is small, and we can directly calculate its singular value decomposition. In addition, note that $A$ is fixed in the algorithm, so the SVD of $A$ can be pre-computed and stored.
5 Testing and comparison of algorithm performance
Through the previous tests, we have given specific values for the parameters involved in the algorithm. In this section, we will further test the performance of the algorithm. All the tests are computed on Windows 10 system with Memory 16 GB, CPU Intel(R) Core(TM)i7-8500U CPU1.80GHz by MATLAB 2016b. We will compare the boundary interval algorithm (M1) proposed in this paper with Lyons algorithm (M2) and the Implicit projection algorithm (M3) in [14]. In the other two algorithms, we take the parameters as $T=2,\gamma=2$.
Figure 11 shows execution time for increasing the number of nodes $M$ of three algorithms. According to the parameter settings in (28), the algorithm only needs to perform an SVD of a matrix with a dimension of $40\times 41$, an IFFT of length $L_{\Delta}=2\times 6\times 20=240$, and an FFT of length $12M+$ 1. As $M$ increases, the speed of the present algorithm becomes apparent.
Figure 11: Execution time for increasing the number of nodes $M$, for MATLAB implementations of the Boundary Interval algorithm (M1), the Lyon algorithm (M2) and Implicit projection algorithm (M3).
The accuracy of the solution obtained through three algorithms is shown in Fig. 12 through 14. In order to better show the differences, we divide the function into three cases for comparison:
  • Overall smooth functions:
    $f_{1}(t)=t^{2},\quad f_{2}(t)=\exp(t),\quad f_{3}(t)=e^{\sin(2.7\pi t)+\cos(\pi t)}.$
  • Internal oscillating functions:
    $f_{4}(t)=\frac{1}{1+100t^{2}},\quad f_{5}(t)=\frac{100}{\cos(1+25t^{2})},\quad f_{6}(t)=\text{erf}(100t).$
  • Boundary oscillating functions:
    $f_{7}(t)=\frac{1}{1.1-t^{2}},\quad f_{8}(t)=\text{Ai}(-66-70t),\quad f_{9}(t)=e^{\sin(65.5\pi t-27\pi)-\cos(20.6\pi t)}.$
(a) $f(t)=t^{2}$
(b) approximation error
(c) $f(t)=\exp(t)$
(d) approximation error
(e) $f(t)=\exp(\sin(2.7\pi t)+\cos(\pi t))$
(f) approximation error
Figure 12: Overall smooth functions and their approximation errors against $M$.
As can be seen from Figure 12, for the first case, when the function is overall smooth, the convergence of the present algorithm is slightly better than the other two algorithms, and it reaches the ideal accuracy at a smaller $M$. At this time, the number of nodes $M$ required by the three algorithms to achieve machine accuracy is relatively small, so there is no obvious difference in performance.
But as shown in Figure 13, for the second case, that is, when the function oscillates internally, since the method proposed in this article relies more on boundary data, the impact of this singularity on it is much smaller than the other two algorithms. The number of nodes $M$ required to achieve machine accuracy is significantly reduced (about $1/4$ of the other two algorithms).
For the third case, since the function oscillates at the boundary, the algorithm in this paper cannot avoid this singularity, so we see that the convergence performance of the three algorithms is very close. However, it should be noted that the number of nodes $M$ required by the three algorithms to achieve machine accuracy is very large at this time. Therefore, as shown in Figure 3, this method still has an advantage in computational complexity.
(a) $f(t)=\frac{1}{1+100t^{2}}$
(b) approximation error
(c) $f(t)=\cos\left(\frac{100}{1+25t^{2}}\right)$
(d) approximation error
(e) $f(t)=\text{erf}(100t)$
(f) approximation error
Figure 13: Internal oscillating functions and their approximation errors against $M$.
(a) $f(t)=\frac{1}{1.1-t^{2}}$
(b) approximation error
(c) $f(t)=\text{Ai}(-66-70t)$
(d) approximation error
(e) $f(t)=\exp(\sin(65.5\pi t-27\pi)-\cos(20.6\pi t))$
(f) approximation error
Figure 14: Boundary oscillating functions and their approximation errors against $M$.
6 Conclusions
A new fast algorithm based on boundary interval data has been proposed for the computation of the Fourier extension approximation to nonperiodic functions. We tested the relevant parameters and provided the setting scheme. The final algorithm has a computational complexity of $O(M\log(M))$ and can reduce the impact of internal singularities on reconstruction. This algorithm can be used as a basic tool for Fourier extension calculation.

### v4

@@
-A Fast Algorithms for Fourier extension based on boundary interval data
+Fast Algorithms for Fourier extension based on boundary interval data thanks: The research is partially supported by National Natural Science Foundation of China (No. 12171455, RSF-NSFC 23-41-00002)
 Abstract
-This paper present a new algorithm for the computation of Fourier extension based on boundary data, which can obtain a super-algebraic convergent Fourier approximation for non-periodic functions.
-The algorithm calculates the extension part through boundary data and connects it with the original function to form a periodic smooth function.
-By testing the key parameters involved, their impact on the algorithm is clarified and the optimization setting scheme of the parameters is proposed.
-Compared with FFT, the algorithm only needs to increase the computational complexity by a fixed small amount.
+In this paper, we first propose a new algorithm for the computation of Fourier extension based on data near the two endpoints of the interval, which can obtain an effective Fourier series approximation for non-periodic functions.
+The algorithm calculates the extended part using the boundary interval data and then combines it with the original data to form the data of the extended function within a period.
+By testing the key parameters involved, their influence on the algorithm was clarified and an optimization setting scheme for the parameters was proposed.
+Compared with FFT, the algorithm only needs to increase the computational complexity by a small amount.
+Then, an improved algorithm for the boundary oscillation function is proposed.
+By refining the boundary interval grid, the resolution constant of the boundary oscillation function was reduced to approximately $1/4$ of the original method.
 keywords:
-Fourier Series, Fourier Extensions, Fourier Continuation
+Fourier Series, Fourier Extensions, Fourier Continuation, Resolution Constant
 {AMS}
@@
 1 Introduction
-Fourier series approximation for smooth periodic functions is popular because it can offer spectral convergence and be computed numerically via the FFT.
-However, the situation is changed completely once periodicity is lost.
-A jump discontinuity is imposed in the approximation at the domain boundaries due to the well known Gibbs phenomenon.
-Many ways have been developed to ameliorate the ill effects of the Gibbs phenomenon, including the methods based on Gegenbauer polynomials [9, 10, 16], the methods utilizing Padé approximations [7, 8] and the Fourier extension (or continuation) methods [4, 11, 12, 14, 13, 15, 6, 17].
+Fourier series approximation has been widely used in various fields for smooth periodic functions.
+In many physical, engineering and exploration geophysical problems, such as analyzing periodic signals in electrical engineering, studying periodic motion in mechanics such as the oscillation of a pendulum under certain ideal conditions, or decomposing seismic/electromagnetic signals into different frequency components using Fourier series approximation [17, 19, 20].
+The Fourier series approximation offers spectral convergence since more terms are added to the Fourier series, the approximation converges to the actual function in an efficient way in the frequency domain.
+Moreover, it can be computed numerically via the FFT (Fast Fourier Transform), which is a computationally efficient algorithm.
+However, the situation changes completely once the periodicity is lost.
+A jump discontinuity is imposed in the approximation at the domain boundaries, owing to the well-known Gibbs phenomenon.
+For instance, when we attempt to approximate a function that is not periodic over a given interval using Fourier series, we will observe this Gibbs phenomenon near the boundaries of the interval.
+Several methods have been developed to ameliorate the ill effects of the Gibbs phenomenon.
+For example, methods based on Gegenbauer polynomials, which have certain orthogonality properties that can be exploited to obtain better approximate functions [9, 10, 18].
+Methods utilizing Padé approximations, which are rational function approximations, can also be used [7, 8].
+And Fourier extension (or continuation) methods, which extend the function in a certain way to make it more suitable for Fourier-based approximation, are also among the approaches developed to address this problem [4, 6, 11, 12, 13, 14, 15, 21].
 The idea of the existing discrete Fourier extension can be described as follows: Let $f$ be a non-periodic, smooth and sufficiently differentiable function over $[-1,1]$, which is sampled in a predefined set of points $\{t_{\ell}\}\subset[-1,1]$, the discrete Fourier extension $C(f)$ of $f$ is defined by
@@
 $\min\sum_{\ell}\left|f(t_{\ell})-\sum_{k=-N}^{N}{\bf c}_{k}e^{\text{i}\frac{k\pi}{T}t_{\ell}}\right|^{2}.$ (2)
-To obtain a solution to the minimization problem (2), a severely ill-conditioned least-squares system need to be solved.
-In [4], Boyd has introduced the truncated singular value decomposition (TSVD) to stabilize the calculation process, the accuracy almost up to machine precision.
-A convergence analysis by Adcock et.
-al states that for analytic functions the method is at least superalgebraically convergent [2].
-Utilizing randomized algorithms, Lyon in [12] provided a fast algorithm for the computation with $T=2$.
-And in [14], Matthysen et.
-al presented fast algorithms for arbitrary $T$, exploiting it as a specific variant of bandlimited extrapolation.
-This idea was further developed in the literature [15] to form a fast algorithm for function approximation on 2D domains.
-It should be noted that the goal of the extension process is to effectively connect the data at both ends of the function. We construct the following two functions:
-$f_{1}(t)=\left\{\begin{aligned} &-1,&x<-\frac{1}{2}\\ &\tanh\left(\tan\pi t\right),&-\frac{1}{2}\leq t\leq\frac{1}{2}\\ &1,&t>\frac{1}{2}\end{aligned}\right.$ (3)
-$f_{2}(t)=\left\{\begin{aligned} &-1,&t<-\frac{1}{2}\\ &\tanh\left(100*\tan\pi t\right),&-\frac{1}{2}\leq t\leq\frac{1}{2}\\ &1,&t>\frac{1}{2}\end{aligned}\right.$
-Since the data of the two functions near the endpoints are the same, the same connection data can be used on the interval $[-T,1)$ and $(1,T]$.
-An illustration is shown in Fig.
+To obtain a solution to the minimization problem (2), a severely ill-conditioned least-squares system has to be solved.
+In [4], Boyd introduced truncated singular value decomposition (TSVD) to stabilize the calculation process, with an accuracy almost up to machine precision.
+A convergence analysis by Adcock et al.
+states that for analytic functions the method is at least superalgebraically convergent [2].
+Utilizing randomized algorithms, Lyon provided a fast algorithm with $T=2$ [12].
+In [14], Matthysen et al.
+presented fast algorithms for arbitrary $T$, exploiting it as a specific variant of bandlimited extrapolation.
+This idea was further developed in [15] to form a fast algorithm for function approximation in 2D domains.
+Figure 1: Periodic extensions of $f_{1}(t)$ and $f_{2}(t)$.
+It should be noted that the goal of the extension process is to effectively connect the data at both ends of the function. The following two functions were constructed:
+$f_{1}(t)=\left\{\begin{aligned} &-1,&x<-\frac{1}{2},\\ &\tanh\left(\tan\pi t\right),&-\frac{1}{2}\leq t\leq\frac{1}{2},\\ &1,&t>\frac{1}{2},\end{aligned}\right.$ (3)
+$f_{2}(t)=\left\{\begin{aligned} &-1,&t<-\frac{1}{2},\\ &\tanh\left(100\tan\pi t\right),&-\frac{1}{2}\leq t\leq\frac{1}{2},\\ &1,&t>\frac{1}{2}.\end{aligned}\right.$
+Because the data of the two functions near the endpoints are the same, the same connection data can be used in the intervals $[-T,1)$ and $(1,T]$.
+An illustration of this process is shown in Fig.
 1.
-This raises an issue: does the data inside the interval really help the calculation of the extension?
-In fact, A very effective method, known as FC-Gram method has been developed earlier, which uses only boundary data (projecting it into a basis of orthogonal polynomials) for extension calculations [5].
+This raises the issue of whether the data inside the interval really help in the calculation of the extension.
+In fact, a very efficient method, known as the FC-Gram method has been developed earlier, which uses only boundary data (projecting it into a basis of orthogonal polynomials) for extension calculations [5].
 Recently, this method has also been extended to the approximation of two-dimensional functions[6].
-The numerical implementation of the FC-gram algorithm requires a high-precision environment and it is not superalgebraic convergent.
-Figure 1: Periodic extensions of $f_{1}(t)$ and $f_{2}(t)$.
-This paper proposes a new fast algorithm for the calculation of Fourier extension, which has the following traits:
-  • Only a few discretization points in the boundary interval are used to calculate the extension part, so the method can be computed with FFT speed.
-Unlike the FC-Gram algorithm, the method in this paper does not need to introduce polynomials.
-  • Since the extension part only uses boundary data for calculation, the influence of the singular points inside the function on the approximation is greatly weakened.
-Therefore, for functions with smooth boundaries and internal oscillations, the proposed algorithm has a smaller re-solution constant than the original full-data algorithm.
-All the results in this paper are based on numerical experiments, and we leave the theoretical proof as a topic for future work.
-The proposed algorithm has good performance and gives clear conclusions on the settings of various parameters.
-Therefore, we feel that it is of substantial interest for those who use FEs in practice.
+In [16], the authors proposed a method to calculate the Fourier extension using the derivative on the boundary combined with a two-point Taylor interpolation polynomial.
+The extension part calculated by these two methods are more stable, but neither of them is superalgebraically convergent because polynomials need to be introduced.
+Judging from the numerical results given in the literature, the number of nodes required by these two methods was significantly higher than that of the previous method when obtaining the same accuracy.
+In [5], the authors identified a critical limitation in directly applying the original extension algorithm with relatively few nodes in the boundary interval: the inherent difficulty in achieving satisfactory approximation accuracy.
+Therefore, they turned to precede the FC(SVD) calculation by projecting them into the bases of orthogonal polynomials.
+However, existing theoretical analyses reveal that the accuracy of the Fourier extension is co-determined by two interdependent parameters –extension length and sampling ratio.
+In previous studies, researchers tended to increase the sampling ratio to improve accuracy.
+This is because in the full data algorithm, increasing the extension length does not seem to bring any benefits.
+In this study, we will point out that by increasing the extension length, the number of nodes required in the boundary interval can be effectively reduced.
+Specifically, we demonstrate that through optimal parameter combinations, an accurate extension computation can be achieved with a few boundary nodes deployment.
+The number of nodes required in the boundary interval can be fixed, and is only related to the desired accuracy.
+In this paper, we first propose a new fast algorithm for the calculation of the Fourier extension, which has the following traits:
+  • By optimizing the parameter settings, only a few nodes in the boundary interval are required to calculate the extension part; Thus, the method can be computed with FFT speed.
+Unlike the FC-Gram algorithm, the method does not require the introduction of polynomials.
+When the behavior of the boundary is not significantly better than that of the interior, the proposed algorithm requires significantly fewer nodes than the FC-Gram algorithm because it maintains the superalgebraic convergence property of the original algorithm.
+  • Because the calculation of the extension part uses only the data of the small interval near the boundary, the influence of oscillations within the interval on the approximation is significantly weakened.
+Therefore, for a function that is smooth in the neighborhood of the interval boundary and oscillates within the interval, the proposed algorithm has a smaller resolution constant than that of the original full-data algorithm.
+This also applies to the FC-Gram algorithm.
+Furthermore, for functions that oscillate in the neighborhood of the interval boundary, we propose an improved algorithm based on boundary interval grid refinement, which can significantly reduce the resolution constant of this type of function.
+The proposed algorithms exhibit good performance and provide clear conclusions regarding the settings of various parameters.
+Therefore, we believe that it is of substantial interest to those who use FEs in practice.
 1.1 Overview of the paper
 In Section 2, we outline the algorithm and explain its structure.
-The specific calculation method of the extension part based on boundary data is given in Section 3.
-Section 4 systematically tests the key parameters, points out their influencing factors, and gives the optimization setting scheme.
-In Section 5, we give the numerical test of the final algorithm and compare it with two other fast algorithms.
-2 The outline of the boundary interval algorithm
+The specific calculation method of the extension part based on boundary interval data is presented in Section 3.
+Section 4 present the computational complexity and resolution constant analysis of the algorithm.
+In Section 5, we systematically tests the key parameters, identifies their influencing factors, and provide an optimization setting scheme.
+In Section 6, we present numerical tests of the final algorithm and compare it with three other fast algorithms.
+Furthermore, for the boundary oscillation function, we propose an improved algorithm based on the boundary interval grid refinement.
+2 Outline of the boundary interval algorithm
 In this paper, we restrict ourselves to the case where the sampling points are uniform, that is,
 $t_{\ell}=\frac{\ell}{M},\quad\ell=-M,\ldots,M.$ (4)
-Let the positive integers $m_{\Delta}$ be the number of nodes taken in the boundary interval, and $S_{l},S_{r}$ be the set of corresponding subscripts,
+Let the positive integer $m_{\Delta}$ be the number of nodes taken in the boundary interval, and $S_{l},S_{r}$ be the set of corresponding subscripts,
 $S_{l}=\{-M,-M+1,\ldots,-M+m_{\Delta}-1\},\quad S_{r}=\{M-m_{\Delta}+1,M-m_{\Delta}+2,\ldots,M\}.$ (5)
-Let $T_{\Delta}$ be a positive interger,
-$L_{\Delta}=2\times T_{\Delta}\times m_{\Delta},\quad h=\frac{2\pi}{L_{\Delta}},\quad x_{j}=(j-1)h,\quad j=1,2,\ldots,L_{\Delta}.$ (6)
-Then let
+Let $T_{\Delta}>1$ and
+$L_{\Delta}=2\times\left\lceil T_{\Delta}\times(m_{\Delta}-1)\right\rceil,\quad h=\frac{2\pi}{L_{\Delta}},\quad x_{j}=(j-1)h,\quad j=1,2,\ldots,L_{\Delta},$ (6)
+[…]
 $J_{\Delta,1}=\{1,2,\ldots,m_{\Delta}\},~{}J_{\Delta,2}=\{\frac{L_{\Delta}}{2}+1,\frac{L_{\Delta}}{2}+2,\ldots,\frac{L_{\Delta}}{2}+m_{\Delta}\},~{}J_{\Delta}=J_{\Delta,1}\cup J_{\Delta,2}.$ (7)
-For $j\in J_{\Delta}$, take
-$g(x_{j})=\left\{\begin{aligned} &f(t_{M-m_{\Delta}+j}),&j\in J_{\Delta,1},\\ &f(t_{-M+j-\frac{L}{2}-1}),&j\in J_{\Delta,2}.\end{aligned}\right.$ (8)
-Then we calculate $g_{c}(x_{j}),j=1,2,\ldots,L_{\Delta}$ according to the method in section 3. Finally, we will obtain a function $f_{c}$ with a period of $2+\lambda$, where
-$\lambda=\frac{L_{\Delta}-2m_{\Delta}+2}{2M},$ (9)
-and the data of the extension function $f_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\frac{L_{\Delta}}{2}-m_{\Delta}$ in one period $[-1,1+\lambda)$ is
-$f_{c}(t_{\ell})=\left\{\begin{aligned} &f(t_{\ell}),&|\ell|\leq M,\\ &g_{c}(x_{m_{\Delta}+\ell-M+1}),&\ell>M.\end{aligned}\right.$ (10)
-Fig.
-2 shows the basic process of the boundary interval algorithm using the function $f(x)=x^{2}$ as an example.
-Here, in order to make the graph look clearer, we use a relatively small number of nodes.
-Figure 1 (a) shows the given data (indicated by the black dotted line) and the boundary data selected for the calculation extension (indicated by the red circle).
-Figure 1 (b) shows the $g_{c}(x_{j})$ calculated from the boundary data.
-The data represented by the blue circle is the subsequent extension data required.
-Figure 1 (c) shows the data of $f_{c}$ in one period formed by combining the original data with the calculated extension data.
-It can be seen from the calculation process that the calculation of the extension part only depends on the boundary interval data.
 (a) Given data and selected boundary data
@@
 Figure 2: Calculation process of a periodic extension of $f(x)=x^{2}$.
-After obtaining the data in one period, the subsequent algorithm can be implemented through FFT.
+For $j\in J_{\Delta}$, take
+$g(x_{j})=\left\{\begin{aligned} &f(t_{M-m_{\Delta}+j}),&j\in J_{\Delta,1},\\ &f(t_{-M+j-{L_{\Delta}}/{2}-1}),&j\in J_{\Delta,2}.\end{aligned}\right.$ (8)
+Then we calculate the extension function $g_{c}$ defined on the interval $[0,2\pi]$ of $g$, and obtain its values at $x_{j},j=1,2,\ldots,L_{\Delta}$ according to the method described in Section LABEL:sec31.
+Finally, we obtain the function $f_{c}$ with a period of $2+\lambda$, where
+$\lambda=\frac{\lceil T_{\Delta}-1\rceil\times(m_{\Delta}-1)}{M},$ (9)
+and the data of the extension function $f_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\frac{L_{\Delta}}{2}-m_{\Delta}$ in one period $[-1,1+\lambda)$ is
+$f_{c}(t_{\ell})=\left\{\begin{aligned} &f(t_{\ell}),&|\ell|\leq M,\\ &g_{c}(x_{m_{\Delta}+\ell-M}),&\ell>M.\end{aligned}\right.$ (10)
+Fig.
+2 shows the basic process of the boundary interval algorithm using the function $f(x)=x^{2}$ as an example.
+Here, to make the graph clearer, we use a relatively small number of nodes.
+Fig.
+2(a) shows the given data (indicated by the black dotted line) and boundary data selected for the calculation extension (indicated by the red circle).
+Fig.
+2(b) shows the $g_{c}(x_{j})$ calculated from the boundary data.
+The data represented by blue circles represent the subsequent extension data required.
+Fig.
+2(c) shows the data of $f_{c}$ in one period formed by combining the original data with the calculated extension data.
+It can be observed from the calculation process that the calculation of the extension part depends only on the boundary interval data.
+After obtaining the data in one period, the subsequent algorithm can be implemented using FFT.
 Therefore, whether the algorithm can achieve satisfactory efficiency depends on the computational complexity of the extension part, that is, the size of the parameters $T_{\Delta}$ and $m_{\Delta}$.
-We will give their selection schemes through testing in Section 4.
+We present their selection schemes through testing in Section 5.
 3 The calculation of $g_{c}$
-Including appropriate normalisation, we let
-$\phi_{k}(x)=\frac{1}{\sqrt{L_{\Delta}}}e^{\text{i}kx},$ (11)
-Assume $g(x_{j})$ is given by (8), then the Fourier extension $g_{c}(x)$ is defined to be
+Including the appropriate normalization, we let
+$\phi_{k}(x)=\frac{1}{\sqrt{L_{\Delta}}}e^{\text{i}kx},\quad k=-n_{\Delta},\ldots,n_{\Delta},\quad x\in[0,2\pi].$ (11)
+Assuming $g(x_{j})$ is given by (8), then the extension function $g_{c}$ of $g$ is defined as
 $g_{c}(x)=\sum_{k=-n_{\Delta}}^{n_{\Delta}}{\bf c}_{k}\phi_{k}(x).$ (12)
@@
 $l=1,2,\ldots,2m_{\Delta},k=-n_{\Delta},\ldots,n_{\Delta}.$
+(a) $n_{\Delta}=100$, $T_{\Delta}=6$
+(b) $m_{\Delta}=100$, $T_{\Delta}=20$
+(c) $m_{\Delta}=100$, $n_{\Delta}=100$
+Figure 3: The singular values of $A$ for various $m_{\Delta}$, $n_{\Delta}$ and $T_{\Delta}$.
 Then the solution of (13) can be obtained by solving the rectangular system
@@
 while ${\bf u}_{i}$ and ${\bf v}_{i}$ are the left and right singular vectors of $A$, respectively. We define
-$\gamma=\frac{m_{\Delta}}{n_{\Delta}}.$ (19)
+$\gamma_{\Delta}=\frac{m_{\Delta}-1}{n_{\Delta}}.$ (19)
 Fig.
 3 shows the singular values of $A$ calculated by SVD techniques for various $m_{\Delta}$, $n_{\Delta}$ and $T_{\Delta}$ in MATLAB.
-It can be seen that the first approximately $\mathcal{O}(n_{\Delta}/T_{\Delta})$ are close to $1$ and then the singular values rapidly decay until the true singular values are masked by machine precision errors.
-(a) $n_{\Delta}=100$, $T_{\Delta}=6$
-(b) $m_{\Delta}=100$, $T_{\Delta}=20$
-(c) $m_{\Delta}=100$, $n_{\Delta}=100$
-Figure 3: The singular values of $A$ for various $m_{\Delta}$, $n_{\Delta}$ and $T_{\Delta}$.
-The situation of singular values reflects the ill-posedness of (15), regularization needs to be introduced.
-Referring to the treatment methods in [2], the truncated SVD leas-squares solution to (15) can be given as
+It can be observed that the first approximately $\mathcal{O}(n_{\Delta}/T_{\Delta})$ are close to $1$ and then the singular values rapidly decay until the true singular values are masked by machine precision errors.
+The situation of singular values reflects the ill-posedness of (15), regularization must be introduced.
+Referring to the treatment methods in [2], the truncated SVD least-squares solution to (15) can be given as
 ${\bf c}_{\tau}=\sum_{\sigma_{i}>\tau}\frac{{\bf u}_{i}^{T}b}{\sigma_{i}}{\bf v}_{i},$ (20)
-where the parameter $\tau$ is usually chosen close to the machine precision.
+where parameter $\tau$ is usually chosen to be close to the machine precision.
+4 Computational complexity and resolution constant
+The computational complexity of the proposed algorithm comprises three parts:
+  • Calculation of $c_{\tau}$.
+The computational effort in this step is mainly concentrated on the singular value decomposition of matrix $A$, which is related to $m_{\Delta}$ and $\gamma$, and can be estimated as
+    • If $\gamma_{\Delta}<1$, the computational complexity is
+      $\mathcal{O}\left(m_{\Delta}^{2}n_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma_{\Delta}}m_{\Delta}^{3}\right).$ (21)
+    • If $\gamma\geq 1$, the computational complexity is
+      $\mathcal{O}\left(m_{\Delta}n^{2}_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma_{\Delta}^{2}}m_{\Delta}^{3}\right).$ (22)
+  • Calculation of $g_{c}$. This step can be achieved by IFFT, the computational complexity is
+    $O\left(L_{\Delta}\log(L_{\Delta})\right).$ (23)
+  • Calculation of Fourier coefficients of $f_{c}$. This can be achieved by FFT, and the computational complexity is
+    $O\left(M\log(M)\right).$ (24)
+Next, we analyze the factors that affect these parameters in detail. First, we provide the following concept for the resolution constant, which comes from the literature [3]:
+Definition 4.1.
+Let $\left\{F_{M}\right\}_{M\in\mathbb{N}}$ be a sequence of approximations such that $F_{M}(f)$ depends only on the values of $f$ on an equispaced grid of $12M+$ 1 points. Let
+$\mathcal{R}(\omega,\delta)=\min\left\{M\in\mathbb{N}:\|\exp(\text{i}\pi\omega t)-F_{M}(\exp(\text{i}\pi\omega t))\|_{\infty}\leq\delta\right\},\quad\omega>0,~{}~{}0<\delta<1,$
+then we say that $F_{M}$ has resolution constant $0<r<\infty$ if
+$R(\omega,\delta)\sim r\omega,\omega\rightarrow\infty,$ (25)
+for any fixed $\delta$.
+The resolution constant of the present method is influenced by two aspects:
+  • First, we need to consider the approximation ability of $g_{c}$ to the boundary interval data of the function $f$.
+For the function $f(x)=\exp(\text{i}\pi\omega t)$ on $[-1,1]$, since the calculation of $g_{c}$ is performed on $[0,2\pi]$ after intercepting the boundary interval data, the function corresponding to $g$ in (8) is
+    $g(x)=\exp(\text{i}\omega_{\Delta}\pi t),$
+    where
+    $\omega_{\Delta}=\frac{m_{\Delta}-1}{M}T_{\Delta}\omega.$ (26)
+    Obviously, if $g_{c}$ is to be able to effectively approximate $g$, it must satisfy
+    $n_{\Delta}\geq\omega_{\Delta}.$ (27)
+    Therefore, according to (26), (27) and (19), it can be obtained that the number of nodes $M$ have to satisfy
+    $M\geq\gamma_{\Delta}T_{\Delta}\omega.$ (28)
+  • Second, it is necessary to consider the number of nodes required for using $f_{c}$ to calculate the overall approximation.
+In this step we use FFT to approximate a periodic function, according to the basic conclusion of Fourier approximation, this step requires the number of nodes $M$ to satisfy
+    $M\geq\omega.$ (29)
+Note that in the above analysis process, we consider the case where the internal and boundary frequencies are the same.
+In the actual calculation process, when facing general functions, it can be seen from the above analysis that if the frequency of the function at the boundary is significantly lower than that of the interior, the number of nodes required for calculation will be greatly reduced.
+This also applies to the FC-Gram algorithm.
+We will verify this in the subsequent numerical experiments.
+Through (4.1)-(4.3) and the above analysis, we can see that whether the proposed method can have a smaller computational complexity depends on whether there is a smaller $m_{\Delta}$ value that can enable $g_{c}$ to accurately approximate the function in the boundary interval.
+In the original full-data Fourier extension algorithm, most literatures recommend the use of oversampling mode.
+The results show that the accuracy of the calculation can be improved with the increase of the sampling ratio
+$\gamma=\frac{M}{N}.$
+In fact, the increase of $\gamma$ plays a role in reducing the discrete step length on $[-T,T]$, which can also be achieved by increasing $T$, as mentioned in the previous literature on the estimation of the resolution constant of the full-data algorithm [3]:
+$r\sim T\gamma.$ (30)
+In previous studies on full-data algorithms, researcher tend to increase $\gamma$ because no benefits were observed with increasing $T$.
+However, for the algorithm proposed in this paper, increasing $T$ seems to be a better choice because it significantly reduces the number of nodes required for the boundary interval.
+In the next section, we will test these parameters more comprehensively and provide clear results and parameter setting suggestions, so that the algorithm can complete the calculation with lower computational complexity.
 5 Numerical investigation and determination of key parameters
-The proposed algorithm involves multiple parameters, and their selection has a substantial impact on the performance of the algorithm.
-This section will test and analyze them and then provide an optimized scheme.
+The proposed algorithm involves multiple parameters, and their selection has a substantial influence on the performance of the algorithm.
+This section tests and analyzes them and then provides an optimized scheme.
 5.1 Testing the parameter $T_{\Delta}$
-Previous research results show that the extension length has a substantial impact on the convergence rate and the resolution power.
-Generally speaking, a larger extension length can achieve a higher convergence rate but also means a larger resolution constant.
-For details, refer to [11, 1, 3].
-In [1], the authors also point out that the theoretical results may not be witnessed in computations.
-For the proposed algorithm, we are more concerned about the influence of $T_{\Delta}$ on the convergence rate, and whether there is any relationship between it and the number of boundary points $m_{\Delta}$, the ratio coefficient $\gamma$, and the frequency of the approximated function.
-Now we take
-$f(t)=\exp(i\pi\omega t)$
-to test different parameters.
-In Fig 4, we give results for the approximation error against $T_{\Delta}$ with different parameters.
-Here and what follows, the approximation error evaluated as the maximum pointwise error over an equi-sampled grid ten times denser than the one used for construction and the parameter $\tau=1e-14$ (we also tested with $\tau=1e-12,1e-13,1e-15$ and obtained results similar to those shown here.)
+From the analysis in the previous section, we know that the extension length $T_{\Delta}$ has a substantial influence on the convergence rate and the resolution power.
+Generally, a larger extension length can achieve a higher convergence rate but also means a larger resolution constant.
+In this section, we numerically observe the impact of $T_{\Delta}$ on the performance of the algorithm. Now we take
+$f(t)=\exp({\text{i}}\pi\omega t)$
+to test the interaction between $T_{\Delta}$ and the other parameters.
+In Fig 4, we present the results of the approximation error against $T_{\Delta}$ for different parameters.
+Here and what follows, the approximation error is evaluated as the maximum pointwise error over an equi-sampled grid that is ten times denser than that used for construction and the parameter $\tau=1e-14$ (we also tested with $\tau=1e-12,1e-13,1e-15$ and obtained results similar to those shown here.) Three distinct regions are visible in all graphs in Fig.
+4:
+  • Region $I_{1}:=\{T_{\Delta}:1<T_{\Delta}<\hat{T}_{1}\}$, where the error decreases rapidly to the machine accuracy as $T_{\Delta}$ increases.
+  • Region $I_{2}:=\{T_{\Delta}:\hat{T}_{1}<T_{\Delta}<\hat{T}_{2}\}$, where the error is maintained close to the machine accuracy.
+  • Region $I_{3}:=\{T_{\Delta}>\hat{T}_{2}\}$, where the error rebounds, and the increase in $T_{\Delta}$ causes the error to increase instead.
 (a) $\gamma=1$, $M=500$, $m_{\Delta}=100$.
@@
 (f) $\gamma=1/2$, $M=1000$, $\omega=20$.
-Figure 4: The approximation error against $T_{\Delta}$ for various parameters
-Three distinct regions are visible in all figures:
-  • A region $I_{1}:=\{T_{\Delta}:1<T_{\Delta}<\hat{T}_{1}\}$, where the error decreases rapidly to near the machine accuracy as $T_{\Delta}$ increases.
-  • A region $I_{2}:=\{T_{\Delta}:\hat{T}_{1}<T_{\Delta}<\hat{T}_{2}\}$, where the error is maintained close to the machine accuracy.
-  • A region $I_{3}:=\{T_{\Delta}>\hat{T}_{2}\}$, where the error rebounds, and the increase of $T_{\Delta}$ will cause the error to increase instead.
-It can be seen that the value of $\hat{T}_{1}$ is influenced by $\gamma$, but not by $\omega$, $m_{\Delta}$, and $M$, while $\hat{T}_{2}$ is influenced by $\gamma$, $M$, and $\omega$.
-For $\hat{T}_{2}$, we can give an estimate:
-$\hat{T}_{2}\sim\frac{M}{\gamma\omega}.$ (21)
-For several different values of $\gamma$, we give the approximate values of $\hat{T}_{1}$ (We tested the functions of $\omega$ from $1$ to $50$, and took the mean of the corresponding $T_{\Delta}$ when their error was less than $135e-$ 13 for the first time as the estimate of $\hat{T}_{1}$), which are listed in Table 1.
-When the number of nodes $M$ is not large enough, the regions $I_{1}$ and $I_{2}$ are merged into one region.
-At this time, the error reaches the optimal value near $\hat{T}_{2}$.
-This phenomenon can be observed from Fig.
-5.
+Figure 4: The approximation error against $T_{\Delta}$ for various parameters.
 Table 1: The approximation value of $\hat{T}_{1}$ for various $\gamma$.
@@
 $\hat{T}_{1}$ 1.09 1.20 2.3 5.9 12.5 24.6 48.9
+It can be seen that the value of $\hat{T}_{1}$ is influenced by $\gamma_{\Delta}$, but not by $\omega$, $m_{\Delta}$, and $M$, while $\hat{T}_{2}$ is influenced by $\gamma_{\Delta}$, $M$, and $\omega$.
+For several different values of $\gamma_{\Delta}$, we give the approximate values of $\hat{T}_{1}$ (we tested the functions of $\omega$ from $1$ to $50$, and took the mean of the corresponding $T_{\Delta}$ when their error was less than $131e-$ 13 for the first time as the estimate of $\hat{T}_{1}$), which are listed in Table 1.
+For $\hat{T}_{2}$, we can estimate:
+$\hat{T}_{2}\sim\frac{M}{\gamma_{\Delta}\omega}.$ (31)
+This is consistent with the result obtained in equation (28).
 (a) $\gamma=1$, $M=100$, $m_{\Delta}=50$.
@@
 Figure 5: The approximation error against $T_{\Delta}$ for $\hat{T}_{2}<\hat{T}_{1}$.
-What will happen if the $T_{\Delta}$ values selected is less than $\hat{T}_{1}$?
+When the number of nodes $M$ is not sufficiently large, regions $I_{1}$ and $I_{2}$ are merged into a single region.
+At this time, the error reached the optimal value near $\hat{T}_{2}$.
+This phenomenon is shown in Fig.
+5.
+What will happen if the $T_{\Delta}$ value selected is less than $\hat{T}_{1}$?
 Figure 6 shows this.
 We can see that the error will not decrease after reaching a certain threshold, which cannot be improved by increasing the numbers of $M$ and $m_{\Delta}$.
-It should be noted that this situation also applies to the Fourier extension algorithm using full data in previous literature.
-In Figure 7, we show the results of the algorithm using full data.
-It can be seen that when the value of $T$ is smaller than the corresponding $\hat{T}_{1}$, the method cannot achieve machine accuracy.
+It should be noted that this situation also applies to the Fourier extension algorithm using the full data in previous literatures.
+In Fig.
+7, we show the results of the algorithm using full data.
+It can be seen that when $\gamma$ is fixed and the value of $T$ is smaller than the corresponding $\hat{T}_{1}$, the method cannot achieve machine accuracy.
 Similarly, this cannot be improved by increasing the number of $M$.
@@
 Figure 7: The approximation error against $M$ for full data FE with $T<\hat{T}_{1}$.
-From the above test results, we can see that the selection of $T_{\Delta}$ is crucial to the approximation accuracy.
-Only when $T_{\Delta}$ is within the interval $[\hat{T}_{1},\hat{T}_{2}]$, the approximation can obtain satisfactory results.
-The length and position of this interval are affected by the parameter $\gamma$ and the frequency $\omega$ of the approximated function.
+From the test results above, we can see that when $\gamma_{\Delta}$ is fixed, the selection of $T_{\Delta}$ is crucial to the approximation accuracy.
+The approximation can obtain satisfactory results only when $T_{\Delta}$ is within the interval $[\hat{T}_{1},\hat{T}_{2}]$.
+The left endpoint $\hat{T}_{1}$ of the interval is determined by parameter $\gamma_{\Delta}$, whereas the right endpoint $\hat{T}_{2}$ is affected by parameters $\gamma_{\Delta}$, $\omega$ and $M$.
 5.2 Testing the parameter $m_{\Delta}$
-The size of parameter $m_{\Delta}$ determines whether this study is meaningful.
-This subsection tests the influences of other parameters on it.
-First, we take $M>{T}_{\Delta}\gamma\omega$ for observation.
-As can be seen from Fig.
+The size of the parameter $m_{\Delta}$ determines whether this study is meaningful.
+This subsection tests the influence of the other parameters on it.
+First, we take $M>{T}_{\Delta}\gamma_{\Delta}\omega$ for observation.
+As shown in Fig.
 8, as the value of $m_{\Delta}$ increases, the error decreases rapidly and reaches the machine precision after a certain value $\hat{m}_{\Delta}$, while the value of $\hat{m}_{\Delta}$ is independent of $M$ and $\omega$.
+(a) $M=1000$, $\gamma_{\Delta}=2$, $T_{\Delta}=6$.
+(b) $\omega=20$, $\gamma_{\Delta}=2$, $T_{\Delta}=6$.
+Figure 8: The approximation error against $m_{\Delta}$ for various $\omega$ and $M$.
+(a) $M=2000$, $\gamma_{\Delta}=8$, $\omega=20$.
+(b) $M=1000$, $\gamma_{\Delta}=4$, $\omega=20$.
+(c) $M=500$, $\gamma_{\Delta}=2$, $\omega=20$.
+(d) $M=500$, $\gamma_{\Delta}=1$, $\omega=20$.
+(e) $M=500$, $\gamma_{\Delta}=1/2$, $\omega=20$.
+(f) $M=500$, $\gamma_{\Delta}=1/4$, $\omega=20$.
+Figure 9: The approximation error against $m_{\Delta}$ for various $\gamma_{\Delta}$ and $T_{\Delta}$.
 In Fig.
-9, we further show how the error changes with $m_{\Delta}$ for different values of $\gamma$ and $T_{\Delta}$.
+9, we further show how the error changes with $m_{\Delta}$ for different values of $\gamma_{\Delta}$ and $T_{\Delta}$.
 It can be seen that the value of $\hat{m}_{\Delta}$ decreases with the increase of $T_{\Delta}$.
-When $\gamma\leq 1$, the value of $\hat{m}_{\Delta}$ at the corresponding $\hat{T}_{1}$ is very close.
-When $\gamma>1$, the value of $\hat{m}_{\Delta}$ at the corresponding $\hat{T}_{1}$ increases rapidly with the increase of $\gamma$.
+This is consistent with the previous analysis, because the step size $h=\frac{2\pi}{L_{\Delta}}$ in the $g_{c}$ calculation process is determined by the product $(m_{\Delta}-1)T_{\Delta}$.
+When $\gamma_{\Delta}\leq 1$, the value of $\hat{m}_{\Delta}$ at the corresponding $\hat{T}_{1}$ is very close.
+When $\gamma_{\Delta}>1$, the value of $\hat{m}_{\Delta}$ at the corresponding $\hat{T}_{1}$ increases rapidly with the increase of $\gamma_{\Delta}$.
 From the perspective of use, we are more concerned about the $\hat{m}_{\Delta}$ value near $\hat{T}_{1}$.
@@
 Table 2: The approximation values of $\hat{m}_{\Delta}$ for various $\gamma$ and $T_{\Delta}$.
-$\gamma$ $8$ $4$ $2$ $1$ $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
+$\gamma_{\Delta}$ $8$ $4$ $2$ $1$ $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
 $T_{\Delta}$ $1.1$ $1.2$ $2.3$ $6$ $12$ $24$ $50$
-$\hat{m}_{\Delta}$ 1820 480 58 17 18 18 19
-(a) $M=1000$, $\gamma=2$, $T_{\Delta}=6$.
-(b) $\omega=20$, $\gamma=2$, $T_{\Delta}=6$.
-Figure 8: The approximation error against $m_{\Delta}$ for various $\omega$ and $M$.
-(a) $M=2000$, $\gamma=8$, $\omega=20$.
-(b) $M=1000$, $\gamma=4$, $\omega=20$.
-(c) $M=500$, $\gamma=2$, $\omega=20$.
-(d) $M=500$, $\gamma=1$, $\omega=20$.
-(e) $M=500$, $\gamma=1/2$, $\omega=20$.
-(f) $M=500$, $\gamma=1/4$, $\omega=20$.
-Figure 9: The approximation error against $m_{\Delta}$ for various $\gamma$ and $T_{\Delta}$.
+$\hat{m}_{\Delta}$ 1870 490 64 23 25 24 26
 (a) $f(t)=\cos\left(\frac{100}{1+25t^{2}}\right)$
 (b) approximation error
-Figure 10: The function and its approximation errors against $m_{\Delta}$ for $M=500$, $\gamma=1$, $T_{\Delta}=6$.
-It is also worth noting that for functions with internal oscillations and smooth boundaries, an increase in the number of $m_{\Delta}$ will cause the value of $\hat{T}_{2}$ to decrease.
-At this time, if the value of $M$ is not large enough, the approximation result using a larger $m_{\Delta}$ may be worse.
-This can be seen from the approximation results for the function $f(t)=\cos\left(\frac{100}{1+25t^{2}}\right)$ in Figure 10.
-4.3 Computational complexity and parameter determination
-The computational complexity of this algorithm consists of three parts:
-  • Calculation of $c_{\tau}$.
-The computational effort in this step is mainly concentrated on the singular value decomposition of the matrix $A$, which is related to $m_{\Delta}$ and $\gamma$, which can be estimated as:
-    • If $\gamma<1$, the computational complexity is
-      $\mathcal{O}\left(m_{\Delta}^{2}n_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma}m_{\Delta}^{3}\right)$ (22)
-    • If $\gamma\geq 1$, the computational complexity is
-      $\mathcal{O}\left(m_{\Delta}n^{2}_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma^{2}}m_{\Delta}^{3}\right)$ (23)
-  • Calculation of $g_{c}$. This step can be achieved by IFFT, the computational complexity is
-    $O\left(L_{\Delta}\log(L_{\Delta})\right)$ (24)
-  • Calculation of Fourier coefficients of $f_{c}$. It can be achieved by FFT, the computational complexity is
-    $O\left(M\log(M)\right).$ (25)
-Combining the test results in the previous two subsections, we know that if we want to approximate the function with frequency $\omega$ with machine precision, the parameters $m_{\Delta}$ and $T_{\Delta}$ need to satisfy
-$m_{\Delta}\geq\hat{m}_{\Delta},\quad T_{\Delta}\geq\hat{T}_{1}.$ (26)
-Accordingly, combined with formula (21), the parameters $L_{\Delta}$ and $M$ need to satisfy
-$L_{\Delta}\geq\hat{L}_{\Delta}=:2\times\hat{T}_{1}\times\hat{m}_{\Delta},\quad M\geq\hat{M}=:\hat{T}_{1}\gamma\omega.$ (27)
-Referring to the results in Table 1 and Table 2, the value of $\hat{M}$ is very close for different $\gamma$ and the value of $\hat{L}_{\Delta}$ is relatively close when $\gamma=1$ and $2$, but the value of $\hat{L}_{\Delta}$ is much larger when $\gamma$ takes other values.
-Further combined with the value of $\hat{m}_{\Delta}$, it can be seen that the computational complexity of the method is equivalent when $\gamma=1$ and $\gamma=2$, while the computational complexity increases significantly in other cases.
-Furthermore, note that a larger $m_{\Delta}$ means a greater probability of encountering high frequencies in the selected boundary interval.
+Figure 10: The function and its approximation errors against $m_{\Delta}$ for $M=500$, $\gamma_{\Delta}=1$, $T_{\Delta}=6$.
+It is also worth noting that for functions with internal oscillations but are smooth near the boundary, as mentioned in Section 5, an increase in the number of $m_{\Delta}$ will cause the value of $\omega$ in (28) to increase.
+If the value of $M$ is not sufficiently large, the approximation result using a larger $m_{\Delta}$ may be worse.
+This can be seen from the approximation results for the function $f(t)=\cos\left(\frac{100}{1+25t^{2}}\right)$ in Fig.
+10.
+5.3 Parameter determination
+Combining the test results in the previous two subsections, we know that if we want to approximate the function with frequency $\omega$ with machine precision, the parameters $T_{\Delta}$ and $m_{\Delta}$ must satisfy
+$T_{\Delta}\geq\hat{T}_{1},\quad m_{\Delta}\geq\hat{m}_{\Delta}.$ (32)
+Accordingly, combined with (31), the parameters $L_{\Delta}$ and $M$ must satisfy
+$L_{\Delta}\geq\hat{L}_{\Delta}=:2\times\hat{T}_{1}\times\hat{m}_{\Delta},\quad M\geq\hat{M}=:\hat{T}_{1}\gamma_{\Delta}\omega.$ (33)
+Referring to the results in Table 1 and Table 2, the value of $\hat{M}$ is very close for different $\gamma_{\Delta}$ values, and the value of $\hat{L}_{\Delta}$ is relatively close when $\gamma_{\Delta}=1$ and $2$; however, the value of $\hat{L}_{\Delta}$ is much larger when $\gamma_{\Delta}$ takes other values.
+Further combined with the value of $\hat{m}_{\Delta}$, it can be seen that the computational complexity of the method is equivalent when $\gamma_{\Delta}=1$ and $\gamma_{\Delta}=2$, whereas the computational complexity increases significantly in other cases.
+Furthermore, a larger $m_{\Delta}$ indicates a greater probability of encountering high frequencies within the selected boundary interval.
 Therefore, we recommend using the following parameters
-$\gamma=1,T_{\Delta}=6,m_{\Delta}=20,$ (28)
+$\gamma_{\Delta}=1,T_{\Delta}=6,m_{\Delta}=25,$ (34)
 for actual calculations, and parameters
-$\gamma=2,T_{\Delta}=2.3,m_{\Delta}=60,$ (29)
+$\gamma_{\Delta}=2,T_{\Delta}=2.3,m_{\Delta}=65,$ (35)
 can be used as an alternative.
-Under such parameter settings, the size of the matrix $A$ in equation (15) is small, and we can directly calculate its singular value decomposition.
+Under such parameter settings, the size of matrix $A$ in (15) is small, and singular value decomposition can be calculated directly.
 In addition, note that $A$ is fixed in the algorithm, so the SVD of $A$ can be pre-computed and stored.
-5 Testing and comparison of algorithm performance
-Through the previous tests, we have given specific values for the parameters involved in the algorithm.
-In this section, we will further test the performance of the algorithm.
-All the tests are computed on Windows 10 system with Memory 16 GB, CPU Intel(R) Core(TM)i7-8500U CPU1.80GHz by MATLAB 2016b.
-We will compare the boundary interval algorithm (M1) proposed in this paper with Lyons algorithm (M2) and the Implicit projection algorithm (M3) in [14].
-In the other two algorithms, we take the parameters as $T=2,\gamma=2$.
-Figure 11 shows execution time for increasing the number of nodes $M$ of three algorithms.
-According to the parameter settings in (28), the algorithm only needs to perform an SVD of a matrix with a dimension of $40\times 41$, an IFFT of length $L_{\Delta}=2\times 6\times 20=240$, and an FFT of length $12M+$ 1.
-As $M$ increases, the speed of the present algorithm becomes apparent.
-Figure 11: Execution time for increasing the number of nodes $M$, for MATLAB implementations of the Boundary Interval algorithm (M1), the Lyon algorithm (M2) and Implicit projection algorithm (M3).
-The accuracy of the solution obtained through three algorithms is shown in Fig. 12 through 14. In order to better show the differences, we divide the function into three cases for comparison:
-  • Overall smooth functions:
-    $f_{1}(t)=t^{2},\quad f_{2}(t)=\exp(t),\quad f_{3}(t)=e^{\sin(2.7\pi t)+\cos(\pi t)}.$
-  • Internal oscillating functions:
-    $f_{4}(t)=\frac{1}{1+100t^{2}},\quad f_{5}(t)=\frac{100}{\cos(1+25t^{2})},\quad f_{6}(t)=\text{erf}(100t).$
-  • Boundary oscillating functions:
-    $f_{7}(t)=\frac{1}{1.1-t^{2}},\quad f_{8}(t)=\text{Ai}(-66-70t),\quad f_{9}(t)=e^{\sin(65.5\pi t-27\pi)-\cos(20.6\pi t)}.$
-(a) $f(t)=t^{2}$
+6 Testing and further improvement of the algorithm
+6.1 Testing and comparison of algorithm performance
+Figure 11: Execution time for increasing the number of nodes $M$, for MATLAB implementations of the Boundary Interval algorithm (M1), the FC-Gram algorithm (M2), the Lyon algorithm (M3) and Implicit projection algorithm (M4).
+Based on the previous tests, we have provided specific values for the parameters involved in the algorithm.
+In this subsection, the performance of the proposed algorithm is tested.
+All tests were computed on a Windows 10 system with 16 GB of memory, Intel(R) Core(TM)i7-8500U CPU1.80GHz using MATLAB 2016b.
+We compared the boundary interval algorithm (M1) proposed in this paper with the FC-Gram algorithm (M2) in [5], the Lyons algorithm (M3) in [12] and the implicit projection algorithm (M4) in [14].
+For convenience of numerical implementation, for M3, we set parameters $T=2,\gamma=2$.
+For M4, we selected the best result from the three parameter combinations of $T=6,\gamma=1$, $T=2,\gamma=2$, and $T=1.2,\gamma=4$ (the performance of the three parameter combinations for different functions is slightly different.
+We also tested other parameter combinations and found that no combination with a product of $T$ and $\gamma$ less than $4$ can achieve machine precision).
+Algorithms M1 and M2 exhibit computational complexity of $\mathcal{O}(M\log M)$ , in contrast to the $\mathcal{O}(M\log^{2}M)$ complexity inherent to algorithms M3 and M4.
+Figure 11 demonstrates the scaling behavior of execution time versus node count $M$ across all four algorithms.
+These experimental observations align with the above conclusions: M1 and M2 demonstrate nearly identical temporal profiles, as do M3 and M4; for equivalent $M$ values, M1/M2 achieve superior computational efficiency compared with M3/M4.
+It should be noted that $M$ values required by these four algorithms are different.
+For M1/M2, the minimum $M$ is governed by the approximation capacity of the boundary interval and maximum frequency of the approximated function.
+For full-data algorithms such as M3 and M4, $M$ depends on both the function frequency and the critical parameter product $T\gamma$.
+Notably, our empirical studies and existing literature establish $T\gamma=4$ as the minimum allowable value for full-data algorithms to attain machine precision.
+Subthreshold $T\gamma$ values result in accuracy degradation for such methods.
+(a) $f(t)=\text{erf}(2t)$
 (b) approximation error
-(c) $f(t)=\exp(t)$
+(c) $f(t)=\text{Ai}(1+3t)$
 (d) approximation error
@@
 Figure 12: Overall smooth functions and their approximation errors against $M$.
-As can be seen from Figure 12, for the first case, when the function is overall smooth, the convergence of the present algorithm is slightly better than the other two algorithms, and it reaches the ideal accuracy at a smaller $M$.
-At this time, the number of nodes $M$ required by the three algorithms to achieve machine accuracy is relatively small, so there is no obvious difference in performance.
-But as shown in Figure 13, for the second case, that is, when the function oscillates internally, since the method proposed in this article relies more on boundary data, the impact of this singularity on it is much smaller than the other two algorithms.
-The number of nodes $M$ required to achieve machine accuracy is significantly reduced (about $1/4$ of the other two algorithms).
-For the third case, since the function oscillates at the boundary, the algorithm in this paper cannot avoid this singularity, so we see that the convergence performance of the three algorithms is very close.
-However, it should be noted that the number of nodes $M$ required by the three algorithms to achieve machine accuracy is very large at this time.
-Therefore, as shown in Figure 3, this method still has an advantage in computational complexity.
 (a) $f(t)=\frac{1}{1+100t^{2}}$
@@
 Figure 13: Internal oscillating functions and their approximation errors against $M$.
-(a) $f(t)=\frac{1}{1.1-t^{2}}$
+The accuracy of the solution obtained using the four algorithms is shown in Fig. 12 - 14. To better show the differences, we divided the function into three cases for comparison:
+  • Overall smooth functions:
+    $f_{1}(t)=\text{erf}(2t),\quad f_{2}(t)=\text{Ai}(1+3t),\quad f_{3}(t)=e^{\sin(2.7\pi t)+\cos(\pi t)}.$
+  • Internal oscillating functions:
+    $f_{4}(t)=\frac{1}{1+100t^{2}},\quad f_{5}(t)=\frac{100}{\cos(1+25t^{2})},\quad f_{6}(t)=\text{erf}(100t).$
+  • Boundary oscillating functions:
+    $f_{7}(t)=\cos(100t^{2}),\quad f_{8}(t)=\text{Ai}(-66-70t),\quad f_{9}(t)=e^{\sin(65.5\pi t-27\pi)-\cos(20.6\pi t)}.$
+(a) $f(t)=\cos(100t^{2})$
 (b) approximation error
@@
 Figure 14: Boundary oscillating functions and their approximation errors against $M$.
+Fig.
+12 shows the numerical results for the first case.
+As mentioned above, the number of nodes $M$ required for M1 and M2 is determined by the frequency of the boundary interval, but M2 cannot achieve superalgebraic convergence using polynomial approximation, so the value of $M$ required is significantly larger than M1; while M3 and M4 are determined by the maximum frequency of the function in the entire interval, and the number of nodes required is slightly larger than M1.
+At this time, the number of nodes $M$ required by the four algorithms to achieve machine accuracy was relatively small, therefore, there was no obvious difference in performance.
+In the second case, that is, when the function oscillates internally.
+As the frequency of the function in the boundary interval is much lower than that in the interior, the number of nodes required for M1 and M2 is determined by the resolution constant of the standard Fourier transform.
+Note that the resolution constants of M3 and M4 are $T\gamma$, so from (29) we can see that the number of nodes required for M1and M2 is approximately $1/4$ of that for M3 and M4, which is consistent with the results given in Fig.
+13.
+In the third case, because the function oscillates near the boundary, the algorithm in this paper cannot avoid this singularity.
+For functions $f_{7}$ and $f_{8}$, the maximum frequency of the function appears in the boundary interval.
+According to (27), the number of nodes required by the proposed algorithm is $T_{\Delta}\gamma_{\Delta}=6$ times the maximum frequency, whereas the number of nodes required by M3 and M4 is $T\gamma=4$ times the maximum frequency.
+This is consistent with the results presented in Figures 2 and 3.
+For $f_{9}$, because the internal frequency is higher, the number of nodes required by the proposed algorithm is less than that of M3 and M4.
+As for M2, we can see that its approximation ability for the boundary interval data is obviously weaker than that of M1, therefore, more nodes are needed to achieve similar accuracy.
+6.2 Algorithm improvement: boundary grid refinement
+In certain situations, different step sizes may be used for interior and boundary sampling.
+The test results in the previous subsection indicate that the resolution constant is affected by the oscillation of the boundary interval data.
+This naturally raises the question: Can the algorithm performance be improved if a finer grid is used near the boundary than inside?
+In this section, we combine this concept to provide a modified algorithm for the boundary oscillation function.
+Let $R$ be an positive integer and
+${t}^{R}_{\ell}=\frac{\ell}{RM},\quad\ell=-RM,\ldots,RM.$ (36)
+The positive integer $m^{R}_{\Delta}$ denotes the number of nodes selected in the boundary interval.
+Notably, only $2\times m^{R}_{\Delta}$ nodes near the boundary were used in the actual calculation.
+The other nodes can be ignored, but they are listed here for convenience.
+For consistency with the boundary interval taken in the previous algorithm, we let
+${m}^{R}_{\Delta}=R(m_{\Delta}-1)+1$ (37)
+and
+$L_{\Delta}^{R}=2\times\left\lceil T_{\Delta}\times(m^{R}_{\Delta}-1)\right\rceil,h^{R}=\frac{2\pi}{L^{R}_{\Delta}},\quad x_{j}^{R}=(j-1)h,\quad j=1,2,\ldots,L^{R}_{\Delta}.$ (38)
+Similarly, let
+$J^{R}_{\Delta,1}=\{1,2,\ldots,m^{R}_{\Delta}\},~{}J^{R}_{\Delta,2}=\{\frac{L^{R}_{\Delta}}{2}+1,\frac{L^{R}_{\Delta}}{2}+2,\ldots,\frac{L^{R}_{\Delta}}{2}+m^{R}_{\Delta}\},~{}J^{R}_{\Delta}=J^{R}_{\Delta,1}\cup J^{R}_{\Delta,2}.$ (39)
+and for $j\in J^{R}_{\Delta}$, take
+$g^{R}(x^{R}_{j})=\left\{\begin{aligned} &f(t^{R}_{RM-m^{R}_{\Delta}+j}),&j\in J_{\Delta,1},\\ &f(t^{R}_{-RM+j-{L^{R}_{\Delta}}/{2}-1}),&j\in J_{\Delta,2}.\end{aligned}\right.$ (40)
+The extension function $g^{R}_{c}(x^{R}_{j}),j=1,2,\ldots,L^{R}_{\Delta}$ was also obtained according to the method described in section 3.
+And then we obtain a function $f^{R}_{c}$ with a period of $2+\lambda$, where $\lambda$ is same as in (9).
+The data of the extension function $f^{R}_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\lceil T_{\Delta}-1\rceil\times(m_{\Delta}-1)-1$ in one period $[-1,1+\lambda)$ are
+$f^{R}_{c}(t_{\ell})=\left\{\begin{aligned} &f(t_{\ell}),&|\ell|\leq M,\\ &g^{R}_{c}(x_{m^{R}_{\Delta}+(\ell-M)\cdot R}),&\ell>M.\end{aligned}\right.$ (41)
+The original algorithm is equivalent to a special case of the improved algorithm when the parameter $R=1$.
+Figure 15: The approximation error against $R$ for functions $\exp(\text{i}\omega\pi t)$, $M=500$.
+(a) $f(t)=\frac{1}{1.01-t^{2}}$
+(b) approximation error
+(c) $f(t)=\text{Ai}(150t)$
+(d) approximation error
+(e) $f(t)=\sin(1500t^{2})$
+(f) approximation error
+Figure 16: Comparisons of algorithms M1 and M4.
+Now we take $\gamma=1,T_{\Delta}=6,m_{\Delta}=25$ to test the influence of parameter $R$ on the algorithm.
+As can be seen from Fig.
+15, when $R\geq 5$, the approximation error not only does not improve, but becomes worse, which is caused by the instability of the calculation of the extension part.
+When $R\leq 4$, it can be observed that the refinement of grid near the boundary improves the approximation error.
+When $\omega=50$, the improvement is not obvious, because the number of nodes $M=500>T_{\Delta}\gamma_{\Delta}\omega$ has exceeded the threshold for resolving this frequency function.
+Therefore, when facing high-frequency functions, we recommend using parameters
+$\gamma_{\Delta}=1,T_{\Delta}=6,m_{\Delta}=25,R=4$
+for calculation.
+Next, we tested the performance of the improved algorithm (M5) and compared it with the original algorithm (M1).
+We selected several test functions with significant high-frequency features in the boundary interval:
+$f_{10}(t)=\frac{1}{1.01-t^{2}},$ $f_{11}(t)=\text{Ai}(150t),$ $f_{12}(t)=\sin(1500t^{2}).$
+Fig.
+16 shows the graphes of the test functions and the corresponding approximation results.
+We can see that the refinement of grid near the boundary effectively reduces the resolution constant.
+The number of nodes required to achieve machine accuracy in the improved method was approximately $1/4$ of that in the original method.
 7 Conclusions
-A new fast algorithm based on boundary interval data has been proposed for the computation of the Fourier extension approximation to nonperiodic functions.
+Fast algorithms based on boundary interval data has been proposed for the computation of the Fourier extension approximation to nonperiodic functions.
 We tested the relevant parameters and provided the setting scheme.
-The final algorithm has a computational complexity of $O(M\log(M))$ and can reduce the impact of internal singularities on reconstruction.
-This algorithm can be used as a basic tool for Fourier extension calculation.
+The new algorithms have a computational complexity of $\mathcal{O}(M\log(M))$ and can be used as basic tools for the calculation of the Fourier extension.
+Acknowledgments
+The authors would like to sincerely thank Mark Lyon for generously sharing the MATLAB codes in [5] and [12] to facilitate comparative analysis in this work.

### v5

@@
-Fast Algorithms for Fourier extension based on boundary interval data thanks: The research is partially supported by National Natural Science Foundation of China (No. 12171455, RSF-NSFC 23-41-00002)
+
 Abstract
-In this paper, we first propose a new algorithm for the computation of Fourier extension based on data near the two endpoints of the interval, which can obtain an effective Fourier series approximation for non-periodic functions.
-The algorithm calculates the extended part using the boundary interval data and then combines it with the original data to form the data of the extended function within a period.
-By testing the key parameters involved, their influence on the algorithm was clarified and an optimization setting scheme for the parameters was proposed.
-Compared with FFT, the algorithm only needs to increase the computational complexity by a small amount.
-Then, an improved algorithm for the boundary oscillation function is proposed.
-By refining the boundary interval grid, the resolution constant of the boundary oscillation function was reduced to approximately $1/4$ of the original method.
-keywords:
-Fourier Series, Fourier Extensions, Fourier Continuation, Resolution Constant
-{AMS}
-42A10, 65T40, 65T50
+This paper presents a novel boundary-optimized fast Fourier extension algorithm for efficient approximation of non-periodic functions.
+The proposed methodology constructs periodic extensions through strategic utilization of boundary interval data, which is subsequently combined with original function samples to form an extended periodic representation.
+We develop a parameter optimization framework that preserves superalgebraic convergence while requiring only a few boundary node deployment, resulting in computational complexity marginally exceeding that of standard FFT implementations.
+Furthermore, we present an improved version of the algorithm tailored for functions exhibiting boundary oscillations.
+This variant employs grid refinement near the boundaries, which reduces the resolution constant to approximately one-fourth of that in conventional approaches.
+Comprehensive numerical experiments confirm the efficiency and accuracy of the proposed methods and establish practical guidelines for parameter selection.
+MSC:
+42A10 41A17 65T40 42C15
 1 Introduction
 Fourier series approximation has been widely used in various fields for smooth periodic functions.
-In many physical, engineering and exploration geophysical problems, such as analyzing periodic signals in electrical engineering, studying periodic motion in mechanics such as the oscillation of a pendulum under certain ideal conditions, or decomposing seismic/electromagnetic signals into different frequency components using Fourier series approximation [17, 19, 20].
+In many physical, engineering and exploration geophysical problems, such as analyzing periodic signals in electrical engineering, studying periodic motion in mechanics such as the oscillation of a pendulum under certain ideal conditions, or decomposing seismic/electromagnetic signals into different frequency components using Fourier series approximation SteinFA2003 ; Tikhonov1995 ; WangIP2011 .
 The Fourier series approximation offers spectral convergence since more terms are added to the Fourier series, the approximation converges to the actual function in an efficient way in the frequency domain.
@@
 Several methods have been developed to ameliorate the ill effects of the Gibbs phenomenon.
-For example, methods based on Gegenbauer polynomials, which have certain orthogonality properties that can be exploited to obtain better approximate functions [9, 10, 18].
-Methods utilizing Padé approximations, which are rational function approximations, can also be used [7, 8].
-And Fourier extension (or continuation) methods, which extend the function in a certain way to make it more suitable for Fourier-based approximation, are also among the approaches developed to address this problem [4, 6, 11, 12, 13, 14, 15, 21].
-The idea of the existing discrete Fourier extension can be described as follows: Let $f$ be a non-periodic, smooth and sufficiently differentiable function over $[-1,1]$, which is sampled in a predefined set of points $\{t_{\ell}\}\subset[-1,1]$, the discrete Fourier extension $C(f)$ of $f$ is defined by
+For example, methods based on Gegenbauer polynomials, which have certain orthogonality properties that can be exploited to obtain better approximate functions 2006Robust ; 1992On ; 2006Optimal .
+Methods utilizing Padé approximations, which are rational function approximations, can also be used 2000A ; Geer1995 .
+And Fourier extension (or continuation) methods, which extend the function in a certain way to make it more suitable for Fourier-based approximation, are also among the approaches developed to address this problem Boyd2002A ; Bruno2022 ; Huybrechs2010 ; Lyon2011 ; 2012Sobolev ; Matthysen2016FAST ; Matthysen2018 ; Webb2020 .
+The idea of the existing discrete Fourier extension can be described as follows: Let $f$ be a non-periodic, smooth and sufficiently differentiable function over [1,1][-1,1], which is sampled in a predefined set of points $\{t_{\ell}\}\subset[-1,1]$, the discrete Fourier extension $C(f)$ of $f$ is defined by
 $C(f)=\sum_{k=-N}^{N}{\bf c}_{k}e^{\text{i}\frac{k\pi}{T}t},\quad T>1,$ (1)
@@
 To obtain a solution to the minimization problem (2), a severely ill-conditioned least-squares system has to be solved.
-In [4], Boyd introduced truncated singular value decomposition (TSVD) to stabilize the calculation process, with an accuracy almost up to machine precision.
+In Boyd2002A , Boyd introduced truncated singular value decomposition (TSVD) to stabilize the calculation process, with an accuracy almost up to machine precision.
 A convergence analysis by Adcock et al.
-states that for analytic functions the method is at least superalgebraically convergent [2].
-Utilizing randomized algorithms, Lyon provided a fast algorithm with $T=2$ [12].
-In [14], Matthysen et al.
+states that for analytic functions the method is at least superalgebraically convergent adcock2014numerical .
+Utilizing randomized algorithms, Lyon provided a fast algorithm with $T=2$ Lyon2011 .
+In Matthysen2016FAST , Matthysen et al.
 presented fast algorithms for arbitrary $T$, exploiting it as a specific variant of bandlimited extrapolation.
-This idea was further developed in [15] to form a fast algorithm for function approximation in 2D domains.
+This idea was further developed in Matthysen2018 to form a fast algorithm for function approximation in 2D domains.
 Figure 1: Periodic extensions of $f_{1}(t)$ and $f_{2}(t)$.
@@
 $f_{2}(t)=\left\{\begin{aligned} &-1,&t<-\frac{1}{2},\\ &\tanh\left(100\tan\pi t\right),&-\frac{1}{2}\leq t\leq\frac{1}{2},\\ &1,&t>\frac{1}{2}.\end{aligned}\right$.
-Because the data of the two functions near the endpoints are the same, the same connection data can be used in the intervals $[-T,1)$ and $(1,T]$.
+Because the data of the two functions near the endpoints are the same, the same connection data can be used in the intervals [T,1)[-T,1) and $(1,T]$.
 An illustration of this process is shown in Fig.
@@
 This raises the issue of whether the data inside the interval really help in the calculation of the extension.
-In fact, a very efficient method, known as the FC-Gram method has been developed earlier, which uses only boundary data (projecting it into a basis of orthogonal polynomials) for extension calculations [5].
-Recently, this method has also been extended to the approximation of two-dimensional functions[6].
-In [16], the authors proposed a method to calculate the Fourier extension using the derivative on the boundary combined with a two-point Taylor interpolation polynomial.
-The extension part calculated by these two methods are more stable, but neither of them is superalgebraically convergent because polynomials need to be introduced.
-Judging from the numerical results given in the literature, the number of nodes required by these two methods was significantly higher than that of the previous method when obtaining the same accuracy.
-In [5], the authors identified a critical limitation in directly applying the original extension algorithm with relatively few nodes in the boundary interval: the inherent difficulty in achieving satisfactory approximation accuracy.
-Therefore, they turned to precede the FC(SVD) calculation by projecting them into the bases of orthogonal polynomials.
-However, existing theoretical analyses reveal that the accuracy of the Fourier extension is co-determined by two interdependent parameters –extension length and sampling ratio.
-In previous studies, researchers tended to increase the sampling ratio to improve accuracy.
-This is because in the full data algorithm, increasing the extension length does not seem to bring any benefits.
-In this study, we will point out that by increasing the extension length, the number of nodes required in the boundary interval can be effectively reduced.
-Specifically, we demonstrate that through optimal parameter combinations, an accurate extension computation can be achieved with a few boundary nodes deployment.
-The number of nodes required in the boundary interval can be fixed, and is only related to the desired accuracy.
-In this paper, we first propose a new fast algorithm for the calculation of the Fourier extension, which has the following traits:
-  • By optimizing the parameter settings, only a few nodes in the boundary interval are required to calculate the extension part; Thus, the method can be computed with FFT speed.
-Unlike the FC-Gram algorithm, the method does not require the introduction of polynomials.
-When the behavior of the boundary is not significantly better than that of the interior, the proposed algorithm requires significantly fewer nodes than the FC-Gram algorithm because it maintains the superalgebraic convergence property of the original algorithm.
+In fact, a very efficient method, known as the FC-Gram method has been developed earlier, which uses only boundary data (projecting it into a basis of orthogonal polynomials) for extension calculations bruno2010high .
+Since then, a lot of research has been done on further improvement and application of the FC-Gram algorithm 2016An ; 2020A ; Bruno2022 .
+In plonka2018numerical , the authors proposed a method to calculate the Fourier extension using the derivative on the boundary combined with a two-point Taylor interpolation polynomial.
+The extension part calculated by these two methods are more stable, but neither of them is superalgebraically convergent.
+When the boundary interval is more volatile than the interior, the number of nodes required to achieve near-machine precision by these two methods was higher than that of the previous method.
+In bruno2010high , the authors identified a critical limitation in directly applying the original extension algorithm with relatively few nodes in the boundary interval: the inherent difficulty in achieving satisfactory approximation accuracy.
+To address this, they introduced a pre-processing step involving Gram polynomial basis generation.
+Existing theoretical analyses reveal that Fourier extension accuracy is co-determined by two interdependent parameters - extension length and sampling ratio.
+While previous studies prioritized increasing sampling ratios for accuracy improvement (as extension length adjustments offered limited benefits in full-data algorithms), this work demonstrates that strategically increasing extension length can effectively reduce boundary node requirements.
+Specifically, we establish that through optimal parameter combinations, accurate extension computation can be achieved with a few boundary node deployment, where the required node count becomes fixed and solely dependent on target accuracy.
+This paper presents a novel fast Fourier extension algorithm with the following key innovations:
+  • Parameter optimization enables extension computation using minimal boundary nodes while maintaining FFT-compatible efficiency.
+Unlike FC-Gram approaches, our method eliminates polynomial basis requirements.
   • Because the calculation of the extension part uses only the data of the small interval near the boundary, the influence of oscillations within the interval on the approximation is significantly weakened.
@@
 $t_{\ell}=\frac{\ell}{M},\quad\ell=-M,\ldots,M.$ (4)
-Let the positive integer $m_{\Delta}$ be the number of nodes taken in the boundary interval, and $S_{l},S_{r}$ be the set of corresponding subscripts,
-$S_{l}=\{-M,-M+1,\ldots,-M+m_{\Delta}-1\},\quad S_{r}=\{M-m_{\Delta}+1,M-m_{\Delta}+2,\ldots,M\}.$ (5)
-Let $T_{\Delta}>1$ and
-$L_{\Delta}=2\times\left\lceil T_{\Delta}\times(m_{\Delta}-1)\right\rceil,\quad h=\frac{2\pi}{L_{\Delta}},\quad x_{j}=(j-1)h,\quad j=1,2,\ldots,L_{\Delta},$ (6)
-[…]
-$J_{\Delta,1}=\{1,2,\ldots,m_{\Delta}\},~{}J_{\Delta,2}=\{\frac{L_{\Delta}}{2}+1,\frac{L_{\Delta}}{2}+2,\ldots,\frac{L_{\Delta}}{2}+m_{\Delta}\},~{}J_{\Delta}=J_{\Delta,1}\cup J_{\Delta,2}.$ (7)
-(a) Given data and selected boundary data
-(b) $g_{c}(x_{j})$ with $T_{\Delta}=4$
-(c) $f_{c}(t_{\ell})$
-Figure 2: Calculation process of a periodic extension of $f(x)=x^{2}$.
-For $j\in J_{\Delta}$, take
+Let $m_{\Delta}\in\mathbb{N}^{+}$ denote the number of boundary interval nodes, with corresponding index sets
+$S_{l}=\{-M,-M+1,\ldots,-M+m_{\Delta}-1\},$ (5)
+$S_{r}=\{M-m_{\Delta}+1,M-m_{\Delta}+2,\ldots,M\}.$
+The calculation of the extension part is performed in the interval $0,2\pi]$.
+Define $T_{\Delta}>1$ as the scaling ratio between the extended interval $0,2\pi]$ and the data-containing interval, analogous to parameter $T$ in the original full-data algorithm.
+Further determine the number of nodes $L_{\Delta}$ used in the extension calculation and the corresponding discrete node $\{x_{j}\}$, specifically
+$L_{\Delta}=2\times\left\lceil T_{\Delta}\times(m_{\Delta}-1)\right\rceil,$ (6)
+$h=\frac{2\pi}{L_{\Delta}},\quad x_{j}=(j-1)h,\quad j=1,2,\ldots,L_{\Delta},$
+where $\lceil\cdot\rceil$ is the rounding symbol. Then let
+$J_{\Delta,1}=\{1,2,\ldots,m_{\Delta}\},~J_{\Delta,2}=\{\frac{L_{\Delta}}{2}+1,\frac{L_{\Delta}}{2}+2,\ldots,\frac{L_{\Delta}}{2}+m_{\Delta}\},$ (7)
+$J_{\Delta}=J_{\Delta,1}\cup J_{\Delta,2}.$
+They indicate the locations of the boundary data during the extension calculation process. For $j\in J_{\Delta}$, take
 $g(x_{j})=\left\{\begin{aligned} &f(t_{M-m_{\Delta}+j}),&j\in J_{\Delta,1},\\ &f(t_{-M+j-{L_{\Delta}}/{2}-1}),&j\in J_{\Delta,2}.\end{aligned}\right$. (8)
-Then we calculate the extension function $g_{c}$ defined on the interval $[0,2\pi]$ of $g$, and obtain its values at $x_{j},j=1,2,\ldots,L_{\Delta}$ according to the method described in Section LABEL:sec31.
+Then we calculate the extension function $g_{c}$ defined on the interval $0,2\pi]$ of $g$, and obtain its values at $x_{j},j=1,2,\ldots,L_{\Delta}$ according to the method described in Section 3.
 Finally, we obtain the function $f_{c}$ with a period of $2+\lambda$, where
 $\lambda=\frac{\lceil T_{\Delta}-1\rceil\times(m_{\Delta}-1)}{M},$ (9)
-and the data of the extension function $f_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\frac{L_{\Delta}}{2}-m_{\Delta}$ in one period $[-1,1+\lambda)$ is
+and the data of the extension function $f_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\frac{L_{\Delta}}{2}-m_{\Delta}$ in one period $1,1+\lambda)$ is
 $f_{c}(t_{\ell})=\left\{\begin{aligned} &f(t_{\ell}),&|\ell|\leq M,\\ &g_{c}(x_{m_{\Delta}+\ell-M}),&\ell>M.\end{aligned}\right$. (10)
+(a) Given data and selected boundary data
+(b) $g_{c}(x_{j})$
+(c) $f_{c}(t_{\ell})$
+Figure 2: Calculation process of a periodic extension of $f(x)=\exp(x)$ ($T_{\Delta}=4,M=32$).
 Fig.
-2 shows the basic process of the boundary interval algorithm using the function $f(x)=x^{2}$ as an example.
+2 shows the basic process of the boundary interval algorithm using the function $f(x)=\exp(x)$ as an example.
 Here, to make the graph clearer, we use a relatively small number of nodes.
@@
 If we let
-[…]
-[…]
+$A_{l,k}$ $=$ $\left\{\begin{aligned} &\phi_{k}(x_{l}),&1\leq l\leq m_{\Delta},\\ &\phi_{k}(x_{\frac{L_{\Delta}}{2}+l-m_{\Delta}}),&m_{\Delta}<l\leq 2m_{\Delta},\end{aligned}\right$. (14)
+$b_{l}$ $=$ $\left\{\begin{aligned} &\frac{1}{m_{\Delta}}f(x_{l}),&1\leq l\leq m_{\Delta},\\ &\frac{1}{m_{\Delta}}f(x_{\frac{L_{\Delta}}{2}+l-m_{\Delta}}),&m_{\Delta}<l\leq 2m_{\Delta},\end{aligned}\right$.
 $l=1,2,\ldots,2m_{\Delta},k=-n_{\Delta},\ldots,n_{\Delta}.$
@@
 (c) $m_{\Delta}=100$, $n_{\Delta}=100$
-Figure 3: The singular values of $A$ for various $m_{\Delta}$, $n_{\Delta}$ and $T_{\Delta}$.
+Figure 3: The singular values of AA for various $m_{\Delta}$, $n_{\Delta}$ and $T_{\Delta}$.
 Then the solution of (13) can be obtained by solving the rectangular system
@@
 The situation of singular values reflects the ill-posedness of (15), regularization must be introduced.
-Referring to the treatment methods in [2], the truncated SVD least-squares solution to (15) can be given as
+Referring to the treatment methods in adcock2014numerical , the truncated SVD least-squares solution to (15) can be given as
 ${\bf c}_{\tau}=\sum_{\sigma_{i}>\tau}\frac{{\bf u}_{i}^{T}b}{\sigma_{i}}{\bf v}_{i},$ (20)
@@
 4 Computational complexity and resolution constant
-The computational complexity of the proposed algorithm comprises three parts:
+The computational complexity of the proposed algorithm decomposes into three principal components:
   • Calculation of $c_{\tau}$.
-The computational effort in this step is mainly concentrated on the singular value decomposition of matrix $A$, which is related to $m_{\Delta}$ and $\gamma$, and can be estimated as
+The computational effort in this step is mainly concentrated on the singular value decomposition of matrix $A$, which is related to $m_{\Delta}$ and $\gamma_{\Delta}$, and can be estimated as
     • If $\gamma_{\Delta}<1$, the computational complexity is
       $\mathcal{O}\left(m_{\Delta}^{2}n_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma_{\Delta}}m_{\Delta}^{3}\right).$ (21)
-    • If $\gamma\geq 1$, the computational complexity is
+    • If $\gamma_{\Delta}\geq 1$, the computational complexity is
       $\mathcal{O}\left(m_{\Delta}n^{2}_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma_{\Delta}^{2}}m_{\Delta}^{3}\right).$ (22)
@@
     $O\left(M\log(M)\right).$ (24)
-Next, we analyze the factors that affect these parameters in detail. First, we provide the following concept for the resolution constant, which comes from the literature [3]:
-Definition 4.1.
-Let $\left\{F_{M}\right\}_{M\in\mathbb{N}}$ be a sequence of approximations such that $F_{M}(f)$ depends only on the values of $f$ on an equispaced grid of $12M+$ 1 points. Let
-$\mathcal{R}(\omega,\delta)=\min\left\{M\in\mathbb{N}:\|\exp(\text{i}\pi\omega t)-F_{M}(\exp(\text{i}\pi\omega t))\|_{\infty}\leq\delta\right\},\quad\omega>0,~{}~{}0<\delta<1,$
+Next, we analyze the factors that affect these parameters in detail. First, we provide the following concept for the resolution constant, which comes from the literature adcock2014parameter :
+Definition 1
+Let $\left\{F_{M}\right\}_{M\in\mathbb{N}}$ be a sequence of approximations such that $F_{M}(f)$ depends only on the values of $f$ on an equispaced grid of $2M+1$ points.
+For $\omega>0,~~0<\delta<1,$ let
+$\mathcal{R}(\omega,\delta)=\min\left\{M\in\mathbb{N}:\|\exp(\text{i}\pi\omega t)-F_{M}(\exp(\text{i}\pi\omega t))\|_{\infty}\leq\delta\right\},$
 then we say that $F_{M}$ has resolution constant $0<r<\infty$ if
-$R(\omega,\delta)\sim r\omega,\omega\rightarrow\infty,$ (25)
+$R(\omega,\delta)\sim r\omega,\quad\omega\rightarrow\infty,$ (25)
 for any fixed $\delta$.
@@
   • First, we need to consider the approximation ability of $g_{c}$ to the boundary interval data of the function $f$.
-For the function $f(x)=\exp(\text{i}\pi\omega t)$ on $[-1,1]$, since the calculation of $g_{c}$ is performed on $[0,2\pi]$ after intercepting the boundary interval data, the function corresponding to $g$ in (8) is
+For the function $f(x)=\exp(\text{i}\pi\omega t)$ on [1,1][-1,1], since the calculation of $g_{c}$ is performed on $0,2\pi]$ after intercepting the boundary interval data, the function corresponding to $g$ in (8) is
     $g(x)=\exp(\text{i}\omega_{\Delta}\pi t),$
@@
 $\gamma=\frac{M}{N}.$
-In fact, the increase of $\gamma$ plays a role in reducing the discrete step length on $[-T,T]$, which can also be achieved by increasing $T$, as mentioned in the previous literature on the estimation of the resolution constant of the full-data algorithm [3]:
+In fact, the increase of $\gamma$ plays a role in reducing the discrete step length on [T,T][-T,T], which can also be achieved by increasing $T$, as mentioned in the previous literature on the estimation of the resolution constant of the full-data algorithm adcock2014parameter :
 $r\sim T\gamma.$ (30)
@@
   • Region $I_{3}:=\{T_{\Delta}>\hat{T}_{2}\}$, where the error rebounds, and the increase in $T_{\Delta}$ causes the error to increase instead.
-(a) $\gamma=1$, $M=500$, $m_{\Delta}=100$.
-(b) $\gamma=1$, $M=1000$, $\omega=20$.
-(c) $\gamma=2$, $M=500$, $m_{\Delta}=150$.
-(d) $\gamma=2$, $M=1000$, $\omega=20$.
-(e) $\gamma=1/2$, $M=500$, $m_{\Delta}=50$.
-(f) $\gamma=1/2$, $M=1000$, $\omega=20$.
+(a) $\gamma_{\Delta}=1$, $M=500$, $m_{\Delta}=100$.
+(b) $\gamma_{\Delta}=1$, $M=500$, $\omega=20$.
+(c) $M=500$, $m_{\Delta}=100$, $\omega=20$.
+(d) $\gamma_{\Delta}=2$, $m_{\Delta}=100$, $\omega=20$.
 Figure 4: The approximation error against $T_{\Delta}$ for various parameters.
 Table 1: The approximation value of $\hat{T}_{1}$ for various $\gamma$.
-$\gamma$ $8$ $4$ $2$ $1$ $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
+$\gamma$ 88 44 22 11 $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
 $\hat{T}_{1}$ 1.09 1.20 2.3 5.9 12.5 24.6 48.9
 It can be seen that the value of $\hat{T}_{1}$ is influenced by $\gamma_{\Delta}$, but not by $\omega$, $m_{\Delta}$, and $M$, while $\hat{T}_{2}$ is influenced by $\gamma_{\Delta}$, $M$, and $\omega$.
-For several different values of $\gamma_{\Delta}$, we give the approximate values of $\hat{T}_{1}$ (we tested the functions of $\omega$ from $1$ to $50$, and took the mean of the corresponding $T_{\Delta}$ when their error was less than $131e-$ 13 for the first time as the estimate of $\hat{T}_{1}$), which are listed in Table 1.
+For several different values of $\gamma_{\Delta}$, we give the approximate values of $\hat{T}_{1}$ (we tested the functions of $\omega$ from $1$ to $50$, and took the mean of the corresponding $T_{\Delta}$ when their error was less than 1e131e-13 for the first time as the estimate of $\hat{T}_{1}$), which are listed in Table 1.
 For $\hat{T}_{2}$, we can estimate:
@@
 This is consistent with the result obtained in equation (28).
-(a) $\gamma=1$, $M=100$, $m_{\Delta}=50$.
+(a) $\gamma 1=$, $M=100$, $m_{\Delta}=50$.
 (b) $\gamma=2$, $M=400$, $m_{\Delta}=100$.
@@
 From the test results above, we can see that when $\gamma_{\Delta}$ is fixed, the selection of $T_{\Delta}$ is crucial to the approximation accuracy.
-The approximation can obtain satisfactory results only when $T_{\Delta}$ is within the interval $[\hat{T}_{1},\hat{T}_{2}]$.
+The approximation can obtain satisfactory results only when $T_{\Delta}$ is within the interval $\hat{T}_{1},\hat{T}_{2}]$.
 The left endpoint $\hat{T}_{1}$ of the interval is determined by parameter $\gamma_{\Delta}$, whereas the right endpoint $\hat{T}_{2}$ is affected by parameters $\gamma_{\Delta}$, $\omega$ and $M$.
-5.2 Testing the parameter $m_{\Delta}$
-The size of the parameter $m_{\Delta}$ determines whether this study is meaningful.
-This subsection tests the influence of the other parameters on it.
-First, we take $M>{T}_{\Delta}\gamma_{\Delta}\omega$ for observation.
-As shown in Fig.
-8, as the value of $m_{\Delta}$ increases, the error decreases rapidly and reaches the machine precision after a certain value $\hat{m}_{\Delta}$, while the value of $\hat{m}_{\Delta}$ is independent of $M$ and $\omega$.
 (a) $M=1000$, $\gamma_{\Delta}=2$, $T_{\Delta}=6$.
@@
 Figure 9: The approximation error against $m_{\Delta}$ for various $\gamma_{\Delta}$ and $T_{\Delta}$.
+5.2 Testing the parameter $m_{\Delta}$
+The size of the parameter $m_{\Delta}$ determines whether this study is meaningful.
+This subsection tests the influence of the other parameters on it.
+First, we take $M>{T}_{\Delta}\gamma_{\Delta}\omega$ for observation.
+As shown in Fig.
+8, as the value of $m_{\Delta}$ increases, the error decreases rapidly and reaches the machine precision after a certain value $\hat{m}_{\Delta}$, while the value of $\hat{m}_{\Delta}$ is independent of $M$ and $\omega$.
 In Fig.
@@
 Table 2: The approximation values of $\hat{m}_{\Delta}$ for various $\gamma$ and $T_{\Delta}$.
-$\gamma_{\Delta}$ $8$ $4$ $2$ $1$ $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
-$T_{\Delta}$ $1.1$ $1.2$ $2.3$ $6$ $12$ $24$ $50$
+$\gamma_{\Delta}$ 88 44 22 11 $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
+$T_{\Delta}$ $1.1$ $1.2$ $2.3$ 66 $12$ $24$ $50$
 $\hat{m}_{\Delta}$ 1870 490 64 23 25 24 26
@@
 All tests were computed on a Windows 10 system with 16 GB of memory, Intel(R) Core(TM)i7-8500U CPU1.80GHz using MATLAB 2016b.
-We compared the boundary interval algorithm (M1) proposed in this paper with the FC-Gram algorithm (M2) in [5], the Lyons algorithm (M3) in [12] and the implicit projection algorithm (M4) in [14].
-For convenience of numerical implementation, for M3, we set parameters $T=2,\gamma=2$.
-For M4, we selected the best result from the three parameter combinations of $T=6,\gamma=1$, $T=2,\gamma=2$, and $T=1.2,\gamma=4$ (the performance of the three parameter combinations for different functions is slightly different.
-We also tested other parameter combinations and found that no combination with a product of $T$ and $\gamma$ less than $4$ can achieve machine precision).
+We compared the boundary interval algorithm (M1) proposed in this paper with the FC-Gram algorithm (M2) in bruno2010high , the Lyons algorithm (M3) in Lyon2011 and the implicit projection algorithm (M4) in Matthysen2016FAST .
+For M2, we use the codes at https://github.com/oscarbruno/FC and take the parameters $d=13$, $C=33$.
+For the specific meaning of the parameters, please refer to 2016An .
+We also tested other parameter values.
+As $d$ increases, the algorithm becomes unstable.
+A large $d$ value (such as $d=30$) will prevent the algorithm from reaching near machine precision.
+According to our test results, if our approximation goal is to achieve near machine accuracy when $M$ is relatively small, then the optimal value of $d$ is approximately $12\sim 15$.
+For M3, we set parameters $T=2,\gamma=2$.
+For M4, we selected the best result from the three parameter combinations of $T=6,\gamma=1$; $T=2,\gamma=2$ and $T=1.2,\gamma=4$ (the performance of the three parameter combinations for different functions is slightly different.
+We also tested other parameter combinations and found that no combination with a product of $T$ and $\gamma$ less than $4$ can achieve near machine precision).
 Algorithms M1 and M2 exhibit computational complexity of $\mathcal{O}(M\log M)$, in contrast to the $\mathcal{O}(M\log^{2}M)$ complexity inherent to algorithms M3 and M4.
@@
   • Boundary oscillating functions:
-    $f_{7}(t)=\cos(100t^{2}),\quad f_{8}(t)=\text{Ai}(-66-70t),\quad f_{9}(t)=e^{\sin(65.5\pi t-27\pi)-\cos(20.6\pi t)}.$
+[…]
+    $f_{9}(t)=e^{\sin(65.5\pi t-27\pi)-\cos(20.6\pi t)}.$
 (a) $f(t)=\cos(100t^{2})$
@@
 12 shows the numerical results for the first case.
-As mentioned above, the number of nodes $M$ required for M1 and M2 is determined by the frequency of the boundary interval, but M2 cannot achieve superalgebraic convergence using polynomial approximation, so the value of $M$ required is significantly larger than M1; while M3 and M4 are determined by the maximum frequency of the function in the entire interval, and the number of nodes required is slightly larger than M1.
+As mentioned above, the number of nodes $M$ required for M1 and M2 is determined by the frequency of the boundary interval, but the re-solution ability of polynomial approximation is weaker than that of trigonometric approximation, so the value of $M$ required by M2 is larger than M1; while M3 and M4 are determined by the maximum frequency of the function in the entire interval, and the number of nodes required is slightly larger than M1.
 At this time, the number of nodes $M$ required by the four algorithms to achieve machine accuracy was relatively small, therefore, there was no obvious difference in performance.
@@
 For $f_{9}$, because the internal frequency is higher, the number of nodes required by the proposed algorithm is less than that of M3 and M4.
-As for M2, we can see that its approximation ability for the boundary interval data is obviously weaker than that of M1, therefore, more nodes are needed to achieve similar accuracy.
+As for M2, we can see that its approximation ability for the boundary interval data is obviously weaker than that of M1, therefore, more nodes are needed to achieve near-machine accuracy.
+We observe that in certain cases (e.g., Fig.
+12(f)), Algorithm M2 exhibits superior accuracy over M1 for small node counts $M$.
+Since M2 employs fewer boundary nodes, its captured boundary intervals inherently exhibit smoother relative behavior compared to M1 under the same conditions.
+Notably, reducing the number of boundary nodes in M1 can similarly enhance early-stage accuracy; however, as demonstrated in prior tests (Section 5.2), such parameter configurations compromise M1’s ability to achieve near-machine precision with sufficient nodes.
 6.2 Algorithm improvement: boundary grid refinement
@@
 and
-$L_{\Delta}^{R}=2\times\left\lceil T_{\Delta}\times(m^{R}_{\Delta}-1)\right\rceil,h^{R}=\frac{2\pi}{L^{R}_{\Delta}},\quad x_{j}^{R}=(j-1)h,\quad j=1,2,\ldots,L^{R}_{\Delta}.$ (38)
+$L_{\Delta}^{R}=2\times\left\lceil T_{\Delta}\times(m^{R}_{\Delta}-1)\right\rceil,$
+$h^{R}=\frac{2\pi}{L^{R}_{\Delta}},\quad x_{j}^{R}=(j-1)h,\quad j=1,2,\ldots,L^{R}_{\Delta}.$
 Similarly, let
-$J^{R}_{\Delta,1}=\{1,2,\ldots,m^{R}_{\Delta}\},~{}J^{R}_{\Delta,2}=\{\frac{L^{R}_{\Delta}}{2}+1,\frac{L^{R}_{\Delta}}{2}+2,\ldots,\frac{L^{R}_{\Delta}}{2}+m^{R}_{\Delta}\},~{}J^{R}_{\Delta}=J^{R}_{\Delta,1}\cup J^{R}_{\Delta,2}.$ (39)
+$J^{R}_{\Delta,1}=\{1,2,\ldots,m^{R}_{\Delta}\},~J^{R}_{\Delta,2}=\{\frac{L^{R}_{\Delta}}{2}+1,\frac{L^{R}_{\Delta}}{2}+2,\ldots,\frac{L^{R}_{\Delta}}{2}+m^{R}_{\Delta}\},$~
+$J^{R}_{\Delta}=J^{R}_{\Delta,1}\cup J^{R}_{\Delta,2}.$
 and for $j\in J^{R}_{\Delta}$, take
@@
 And then we obtain a function $f^{R}_{c}$ with a period of $2+\lambda$, where $\lambda$ is same as in (9).
-The data of the extension function $f^{R}_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\lceil T_{\Delta}-1\rceil\times(m_{\Delta}-1)-1$ in one period $[-1,1+\lambda)$ are
+The data of the extension function $f^{R}_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\lceil T_{\Delta}-1\rceil\times(m_{\Delta}-1)-1$ in one period $1,1+\lambda)$ are
 $f^{R}_{c}(t_{\ell})=\left\{\begin{aligned} &f(t_{\ell}),&|\ell|\leq M,\\ &g^{R}_{c}(x_{m^{R}_{\Delta}+(\ell-M)\cdot R}),&\ell>M.\end{aligned}\right$. (39)
 The original algorithm is equivalent to a special case of the improved algorithm when the parameter $R=1$.
-Figure 15: The approximation error against $R$ for functions $\exp(\text{i}\omega\pi t)$, $M=500$.
+(a) the approximation error against $R$
+(b) the approximation error against $t$ $(\omega=300)$
+Figure 15: The approximation error with various $R$ for functions $\exp(\text{i}\omega\pi t)$, $M=500$.
 (a) $f(t)=\frac{1}{1.01-t^{2}}$
@@
 (f) approximation error
-Figure 16: Comparisons of algorithms M1 and M4.
+Figure 16: Comparisons of algorithms M1 and M5.
 Now we take $\gamma=1,T_{\Delta}=6,m_{\Delta}=25$ to test the influence of parameter $R$ on the algorithm.
 As can be seen from Fig.
-15, when $R\geq 5$, the approximation error not only does not improve, but becomes worse, which is caused by the instability of the calculation of the extension part.
+15(a), when $R\geq 5$, the approximation error not only does not improve, but becomes worse, which is caused by the instability of the calculation of the extension part.
 When $R\leq 4$, it can be observed that the refinement of grid near the boundary improves the approximation error.
@@
 for calculation.
+It can be seen from Fig.
+15(b) demonstrates that progressive increases in parameter $R$ from $1$ to $4$ induce uniform error reduction across the entire computational domain.
+As the boundary grid is refined, the reduction of the boundary error improves the overall approximation error.
 Next, we tested the performance of the improved algorithm (M5) and compared it with the original algorithm (M1).
 We selected several test functions with significant high-frequency features in the boundary interval:
-$f_{10}(t)=\frac{1}{1.01-t^{2}},$ $f_{11}(t)=\text{Ai}(150t),$ $f_{12}(t)=\sin(1500t^{2}).$
+[…]
 Fig.
@@
 The number of nodes required to achieve machine accuracy in the improved method was approximately $1/4$ of that in the original method.
+Next, we compare the performance of the M5 method with Chebyshev interpolation.
+Polynomial interpolation at Chebyshev nodes is a widely used approach for function approximation, renowned for its high accuracy and favorable resolution constant.
+Figure 17 shows the approximation errors for both methods applied to the test functions $f_{3}$, $f_{5}$, $f_{7}$, and $f_{9}$ as the number of interpolation nodes $M$ increases.
+The relative performance of the two methods depends primarily on the characteristics of the function, especially its frequency distribution properties.
+Overall, the performance difference between the methods is modest.
+Furthermore, both techniques are compatible with FFT-based implementations, ensuring high computational efficiency.
+(a) $f_{3}(t)$
+(b) $f_{5}(t)$
+(c) $f_{7}(t)$
+(d) $f_{9}(t)$
+Figure 17: Comparisons of algorithms M5 and the Chebyshev interpolation.
 7 Conclusions
@@
 The new algorithms have a computational complexity of $\mathcal{O}(M\log(M))$ and can be used as basic tools for the calculation of the Fourier extension.
+Conflict of Interest declaration
+The authors declare that they have no affiliations with or involvement in any organization or entity with any financial interest in the subject matter or materials discussed in this manuscript.
 Acknowledgments
-The authors would like to sincerely thank Mark Lyon for generously sharing the MATLAB codes in [5] and [12] to facilitate comparative analysis in this work.
+The authors would like to sincerely thank Mark Lyon for generously sharing the MATLAB codes to facilitate comparative analysis in this work.
+We are also very grateful to the reviewers and the editor for their constructive comments.